### v1

# Ultraluminous X-ray sources in Globular Clusters

11institutetext: Nicolaus Copernicus Astronomical Center, Polish Academy of Sciences, Bartycka 18, 00-716 Warsaw, Poland
22institutetext: Faculty of Mathematics and Computer Science, A. Mickiewicz University, Uniwersytetu Poznańskiego 4, 61-614 Poznań, Poland
Ultraluminous X-ray sources in Globular Clusters
Grzegorz Wiktorowicz E-mail: gwiktoro@camk.edu.pl11    Mirek Giersz 11    Abbas Askar 11    Arkadiusz Hypki 1122    Lucas Helstrom 11
(Accepted XXX. Received YYY; in original form ZZZ)
This paper investigates the formation, populations, and evolutionary paths of UltraLuminous X-ray Sources (ULXs) within Globular Clusters (GCs). ULXs, characterised by their extreme X-ray luminosities, present a challenge to our understanding of accretion physics and compact object formation. While previous studies have largely focused on field populations, this research examines the unique environment of GCs, where dynamical interactions play a significant role. Using the MOCCA Monte Carlo code, we explore how dynamics influences ULX populations within these dense stellar clusters.
Our findings reveal that dynamical processes, such as binary hardening and exchanges, can both facilitate and impede ULX formation in GCs. The study explores the impact of parameters including the initial binary fraction, tidal filling, and multiple stellar populations on the evolution of ULXs. We find that non-tidally filling clusters exhibit significantly larger ULX populations compared to tidally filling ones.
The results indicate that the apparent scarcity of ULXs in GCs may be related to the older stellar populations of GCs relative to the field. Furthermore, the study identifies a population of ”escaper” ULXs, which originate in GCs but are ejected and emit X-rays outside the cluster. These escapers may significantly contribute to the observed field ULX population.
Key Words.:
stellar dynamics – methods: numerical – globular clusters: evolution – X-ray: binaries
1 Introduction
Ultraluminous X-ray sources (ULXs) have been a subject of intense study in the field of high-energy astrophysics for several decades. These enigmatic objects, characterized by X-ray luminosities exceeding $10^{39}\mathrel{\mathrm{erg}/\mathrm{s}}$, challenge our understanding of accretion physics and compact object formation (see e.g. Kaaret et al. 2017a, b, for a recent review). With over $1800$ known ULXs identified to date (Walton et al. 2022), they represent a significant population of extreme X-ray emitters in the Universe, yet their fundamental nature remains a topic of ongoing debate.
The diversity of ULXs has been highlighted in recent classification schemes (e.g. Wiktorowicz et al. 2017), which attempt to categorize these objects based on their observational properties and potential underlying physical mechanisms. While the majority of ULXs were initially thought to harbor intermediate-mass black holes (IMBHs), our understanding has evolved considerably with the detection of pulsar accretors (Bachetti et al. 2014).
The importance of studying ULXs, particularly in the context of globular clusters (GC), cannot be overstated. These objects serve as laboratories for extreme accretion physics, potentially shedding light on the formation and evolution of double compact objects (DCOs) - the progenitors of gravitational wave sources. Moreover, ULXs continue to be considered as candidates for elusive IMBHs (Kaaret et al. 2017a, and references therein).
A critical aspect of ULX research involves identifying and characterizing their donor stars. Heida et al. (2016); López et al. (2020) reported that some ULXs had been identified with potential red super-gian (RSG) donors using infrared spectroscopy data. They examined a total of $113$ ULXs using photometric techniques in the infrared band, resulting in the identification of candidate counterparts for $38$ objects. Among these, the nature of $12$ sources was confirmed through spectroscopic analysis, with only five categorized as ULXs likely to have RSG donors.
In this paper, we present a comprehensive study of ULXs in globular clusters, focusing on their stellar-mass manifestations. The dense stellar environment introduces additional complexities and possibilities for exotic binary formations. By examining their properties, distribution, and potential formation mechanisms, we aim to contribute to the broader understanding of ULX physics and their role in compact object evolution.
1.1 Globular Cluster environment
GCs are dense, quasi-spherical collections of stars that orbit the centers of galaxies. These typically old stellar systems, often containing millions of stars within a relatively small volume, are hubs of dynamical interactions. Their evolution is governed by the relaxation process connected with continuous distance gravitational interactions between all objects in the cluster.
This leads to global phenomena like, core collapse, i.e. contraction of the cluster core and gradual increase in central density, and mass segregation, i.e. the tendency of heavier stars to occupy the central regions, while pushing the lighter stars outwards. Close dynamical interactions influence also the stars and binaries individually, influencing the binary parameters (e.g. the separation and eccentricity), or through exchanges (e.g. binary changes one of it’s component into another star), ejections (star or binary becomes unbound with the cluster), and binary formation (single stars became bound and form a binary). Such dynamical encounters in dense star clusters can also lead to the formation of close binary systems, including X-ray binaries (Pooley et al. 2003; Ivanova et al. 2005, 2008). Understanding how the fundamental dynamics influence the ULX formation is essential for interpreting the presence and characteristics of these exotic objects in GC environments.
Figure 1: A schematic representation of the origins and eventual locations of ULX progenitors and ULXs in GCs. Progenitors can either form and remain bound to the cluster, leading to in-cluster ULX emission (”IN-CLUSTER”), or be dynamically ejected and emit as ULXs outside the cluster (”ESCAPER”). Additionally, ULXs originating and emitting in the field (”FIELD”) represent progenitors formed without direct interaction with cluster dynamics.
Figure 1 illustrates three pathways for ULX formation in relation to GCs. In the ”IN-CLUSTER” scenario, ULX progenitors form and remain bound to the cluster, often through multiple dynamical interactions, including exchanges. ”ESCAPER” ULXs originate in the GC but are ejected through interactions, emitting outside the cluster while retaining their GC-imprinted dynamical history. ”FIELD” ULXs, by contrast, form and evolve in isolation. The key distinction lies in the dynamical imprint: ”ESCAPER” ULXs bear the hallmarks of GC interactions, unlike ”FIELD” systems, which provide a baseline for understanding the unique contribution of clusters to ULX formation.
1.2 GCULX candidates
While the majority of ULXs are detected in star-forming regions of galaxies, there is increasing evidence for their presence in GCs as well. Pioneering observations of GCULXs have been reported in several galaxies, with varied GC environments, spectral behavior, and variability patterns.
Maccarone et al. (2007) discovered the first ULX in a GC (NGC 4472). It exhibits an X-ray luminosity of $4\times 10^{39}\mathrel{\mathrm{erg}/\mathrm{s}}$ and rapid variability. Similar detections were reported in NGC 1399 (Shih et al. 2010; Irwin et al. 2010), with more GCULXs in NGC 4472 (Maccarone et al. 2011) and NGC 4649 (Roberts et al. 2012). Some sources identified in GCs exhibited flares just above the ULX-defining limit of $10^{39}\mathrel{\mathrm{erg}/\mathrm{s}}$ (Irwin et al. 2016; Sivakoff et al. 2005).
Recent observations have extended the known GCULX population, including seven additional GCULXs associated with M87 GCs (Dage et al. 2020), three in NGC 1316 (Dage et al. 2021), and $10$ new candidates in GCs of massive ($>10^{11.5}M_{\odot}$) early-type galaxies (Thygesen et al. 2023).
In dense stellar systems, compact objects can be readily ejected, which can considerably impact the formation channels of such objects (e.g. Kulkarni et al. 1993). Furthermore, GCs typically exhibit high ages and, consequently, fewer higher-mass stars. These characteristics contribute to the intriguing nature of GCULXs.
2 Methodology
2.1 MOCCA Monte Carlo Code
The MOCCA (MOnte Carlo Cluster evolution Code; Hypki & Giersz 2013) is a state-of-the-art numerical tool that combines the Monte Carlo method for stellar dynamics (Hénon 1971; Stodolkiewicz 1986) with detailed stellar and binary evolution algorithms from the BSE code (Hurley et al. 2000, 2002) with recent updates (level-C from Kamlah et al. 2022, and references therein) and low-N scattering code FEWBODY (Fregeau et al. 2004) to follow close dynamical interactions. This hybrid approach makes MOCCA particularly well-suited for studying the formation and evolution of exotic objects like ULXs in dense stellar environments, offering an optimal balance between computational efficiency and physical accuracy.
Recent improvements to the MOCCA code (Hypki et al. 2022, 2025; Giersz et al. 2024) have significantly enhanced its capabilities for ULX studies. First of all, the mass transfer calculations now incorporate a more sophisticated treatment of super-Eddington accretion, crucial for modeling ULX systems (Shakura & Sunyaev 1973; Wiktorowicz et al. 2015). Additionally, improvements in core radii calculations improved the RLOF calculations. The introduction of detailed evolutionary tracking provide unprecedented insight into the spatial distribution and evolutionary pathways of ULX systems within clusters equivalent to detailed outputs from codes like startrack (Belczynski et al. 2008).
A major advancement relevant to this study is MOCCA’s capability to handle multiple stellar populations (MSPs), a phenomenon now recognized as ubiquitous in globular clusters (e.g. Piotto et al. 2015; Renzini et al. 2015; Gratton et al. 2019; Bastian & Lardo 2018; Milone & Marino 2022). While a comprehensive analysis of MSPs in our models is presented in Hypki et al. (2022, 2025); Giersz et al. (2024), here we specifically investigate their impact on ULX formation and evolution. The code now tracks different stellar populations with distinct chemical compositions, ages, and spatial distributions, allowing us to examine how MSP-specific properties affect compact object formation and ULX characteristics.
All recent changes to the MOCCA code are described in details in Giersz et al. (2024).
2.2 Simulations
The MOCCA code employs hundreds of parameters to govern the dynamical and stellar evolution of globular clusters. Based on previous studies of compact object populations - including cataclysmic variables (Belloni et al. 2017), ULXs in the field (Wiktorowicz et al. 2019), and double white dwarf systems (Hellström et al. 2024) - we identified key parameters that most significantly influence ULX formation and evolution. Our parameter selection strategy focused on those that control dynamics and initial stellar population properties. This systematic approach allows us to explore the most relevant parameter space while maintaining computational feasibility. In this section we present discussion of the utilized parameters.
[…]
For all simulations with dynamics enabled, we tracked the evolution of escaped systems to analyze the properties of their populations and potential ULXs that can form after their progenitors are ejected from the cluster. These escapers evolve as isolated binaries, similar to field binary evolution (e.g., Fragos et al. 2015; Wiktorowicz et al. 2019; Zuo et al. 2021). Systems can escape the cluster through multiple mechanisms: (1) dynamical interactions, particularly strong few-body encounters, (2) natal kicks from supernovae explosions, including both direct and Blaauw kicks, (3) gradual two-body relaxation (Spitzer & Hart 1971), and (4) tidal stripping of outer cluster regions by the host galaxy’s gravitational field (Gnedin & Ostriker 1997).
The galactocentric distance ($\mathrm{R}_{\mathrm{gc}}$) for each simulation was scaled to maintain an initial tidal radius of $\mathrm{R}_{\mathrm{tid}}\approx 43\mathrel{\mathrm{pc}}$ across all models. This normalization of the tidal radius facilitates meaningful comparisons between simulations, particularly when analyzing the degree of tidal filling and cluster structural parameters. The tidal radius follows the relation $\mathrm{R}_{\mathrm{tid}}\propto\mathrm{R}_{\mathrm{gc}}^{2/3}\mathrm{M}_{\mathrm{tot}}^{1/3}$ (e.g. Webb et al. 2013), where $\mathrm{M}_{\mathrm{tot}}$ is the cluster mass. This scaling ensures that clusters experience comparable relative tidal forces despite different masses and orbital parameters. We note that while the initial tidal radii are identical, the clusters’ subsequent evolution may lead to different filling factors111filling factor is R_tid/R_h, which is initially $\sim 43$ for non-tidally filling simulations and $\sim 3.6$ for tidally filling ones. depending on their internal dynamics and mass-loss history.
Table 1: Model parameters and initial properties
Label $\mathrm{N}_{\mathrm{pop}}$ $\mathrm{R}_{\mathrm{h,2}}/\mathrm{R}_{\mathrm{h,1}}$ $\mathrm{f}_{\mathrm{bin}}$ $\mathrm{R}_{\mathrm{gc}}$ $\mathrm{R}_{\mathrm{h}}$ $\mathrm{M}_{\mathrm{tot}}/10^{6}$ $\mathrm{R}_{\mathrm{c}}$ $\mathrm{M}_{\mathrm{c}}/10^{3}$ $\log_{10}(\rho_{\mathrm{c}})$ $\mathrm{t}_{\mathrm{h,rel}}$
npop1-fb10-TF 1 10% $2.01$ $11.90$ $0.69$ $5.85$ $89.10$ $2.64$ $10,888.00$
npop1-fb10-nTF 1 10% $2.01$ $1.00$ $0.69$ $0.49$ $89.10$ $5.87$ $265.22$
npop1-fb95-TF 1 95% $1.96$ $11.93$ $0.72$ $5.81$ $93.93$ $2.61$ $6,386.70$
npop1-fb95-nTF 1 95% $1.96$ $1.00$ $0.72$ $0.49$ $93.93$ $5.84$ $155.06$
npop2-cpop05-fb10-TF 2 0.05 10% $2.01$ $11.69$ $0.66$ $0.13$ $12.71$ $356$,505.306505.306,505.306 , 505.30
npop2-cpop05-fb10-TF-NF 2 0.05 10% $2.40$ $11.81$ $0.46$ $6.26$ $68.94$ $2.41$ $9,157.70$
npop2-cpop05-fb10-nTF 2 0.05 10% $2.01$ $0.99$ $0.66$ $0.01$ $12.71$ $9.57$ $160.33$
npop2-cpop05-fb10-nTF-NF 2 0.05 10% $2.40$ $1.00$ $0.46$ $0.53$ $68.94$ $5.63$ $225.73$
npop2-cpop05-fb95-TF 2 0.05 95% $1.96$ $11.75$ $0.69$ $0.12$ $13.03$ $6.37$ $3,790.60$
npop2-cpop05-fb95-TF-NF 2 0.05 95% $2.35$ $11.77$ $0.48$ $6.18$ $71.69$ $2.39$ $5,342.70$
npop2-cpop05-fb95-nTF 2 0.05 95% $1.96$ $1.00$ $0.69$ $0.01$ $13.03$ $9.59$ $93.85$
npop2-cpop05-fb95-nTF-NF 2 0.05 95% $2.35$ $1.00$ $0.48$ $0.53$ $71.69$ $5.61$ $132.31$
npop2-cpop2-fb10-TF 2 0.2 10% $2.01$ $11.80$ $0.66$ $0.50$ $12.84$ $4.55$ $7,083.50$
npop2-cpop2-fb10-TF-NF 2 0.2 10% $2.40$ $11.81$ $0.46$ $6.09$ $64.82$ $2.40$ $9,157.10$
npop2-cpop2-fb10-nTF 2 0.2 10% $2.01$ $1.00$ $0.66$ $0.04$ $12.84$ $7.77$ $174.58$
npop2-cpop2-fb10-nTF-NF 2 0.2 10% $2.40$ $1.00$ $0.46$ $0.52$ $64.82$ $5.61$ $225.72$
npop2-cpop2-fb95-TF 2 0.2 95% $1.96$ $11.77$ $0.69$ $0.49$ $13.12$ $564$,079.504079.504,079.504 , 079.50
npop2-cpop2-fb95-TF-NF 2 0.2 95% $2.35$ $11.77$ $0.48$ $6.12$ $70.32$ $2.35$ $5,342.50$
npop2-cpop2-fb95-nTF 2 0.2 95% $1.96$ $1.00$ $0.69$ $0.04$ $13.12$ $7.77$ $101.01$
npop2-cpop2-fb95-nTF-NF 2 0.2 95% $2.35$ $1.00$ $0.48$ $0.52$ $70.32$ $5.56$ $132.31$
222 List of simulations with differentiating parameters and some initial properties (at $t=0$). Label identifies particular simulation and will be used throughout the paper. TF/nTF in the label means initially tidally filling or non-tidally filling. NF, when present, means that new features were included in this model (see text for detailes); $\mathrm{N}_{\mathrm{pop}}$ number of stellar generations; $\mathrm{R}_{\mathrm{h,2}}/\mathrm{R}_{\mathrm{h,1}}$ ratio between the $\mathrm{R}_{\mathrm{h}}$ of the second and first stellar generation; $\mathrm{f}_{\mathrm{bin}}$ binary fraction; $\mathrm{R}_{\mathrm{gc}}$ galactocentric radius [$\mathrel{\mathrm{kpc}}$]; $\mathrm{R}_{\mathrm{h}}$ half-mass radius for the first population [$\mathrel{\mathrm{pc}}$]; $\mathrm{M}_{\mathrm{tot}}$ total stellar mass [$\mathrel{\mathrm{M}_{\odot}}$]; $\mathrm{R}_{\mathrm{c}}$ core radius [$\mathrel{\mathrm{pc}}$]; $\mathrm{M}_{\mathrm{c}}$ core mass [$\mathrel{\mathrm{M}_{\odot}}$]; $\rho_{\mathrm{c}}$ central density [$\mathrel{\mathrm{M}_{\odot}}/\mathrel{\mathrm{pc}}^{3}$]; $\mathrm{t}_{\mathrm{h,rel}}$ Spitzer half-mass relaxation time [$\mathrel{\mathrm{M}_{\odot}}$]. All simulations include the escapers information (see text). For all these simulations the counterpart field population (with dynamics turned off) was calculated (nodyn in the label; not shown in the table) with the same parameters.
We performed a comprehensive set of simulations exploring different initial conditions and physical parameters. Table 1 summarizes our simulation grid, where each model is identified by a unique label used throughout this paper. The simulations vary in the number of stellar populations ($\mathrm{N}_{\mathrm{pop}}$), with varied spatial distribution of different generations through the concentration parameter ($\mathrm{R}_{\mathrm{h,2}}/\mathrm{R}_{\mathrm{h,1}}$), defined as the ratio between the half-mass radii of subsequent generation relative to the first generation. The initial binary fraction ($\mathrm{f}_{\mathrm{bin}}$) was either $10\%$, or $95\%$, which affected also other parameters (see below). We explored both tidally filling (TF) and non-tidally filling (nTF) configurations ($\mathrm{Tidal}$). Some simulations incorporate the new features (NF) described below, particularly relevant for multiple population scenarios. For each simulation, we provide key structural parameters at $t=0\mathrel{\mathrm{Myr}}$, including the galactocentric radius ($\mathrm{R}_{\mathrm{gc}}$), half-mass radius ($\mathrm{R}_{\mathrm{h}}$), total stellar mass ($\mathrm{M}_{\mathrm{tot}}$), core radius ($\mathrm{R}_{\mathrm{c}}$; according to Casertano & Hut 1985), core mass ($\mathrm{M}_{\mathrm{c}}$, i.e. mass inside $\mathrm{R}_{\mathrm{c}}$), and central density ($\rho_{\mathrm{c}}$; according to Casertano & Hut 1985). Additionally, we computed corresponding field populations with dynamics turned off (denoted by nodyn in the model labels) to serve as control cases, though these are not shown separately in the table.
To better compare simulations with different binary fractions, we made several adjustments to the parameter space. Since a lower binary fraction results in lower total stellar mass (as binary systems typically have higher masses than single stars), we adjusted the initial number of objects to have similar half-mass relaxation time for both models: simulations with $\mathrm{f}_{\mathrm{bin}}=95\%$ used $n=600,000$ objects, while those with $\mathrm{f}_{\mathrm{bin}}=10\%$ used $n=1,063,635$ objects.We also modified the initial distribution of semi-major axes. For simulations with $\mathrm{f}_{\mathrm{bin}}=95\%$, we used the modified version of the Kroupa (1995) period distribution (see Belloni et al. 2017) for stars with $M<5\mathrel{\mathrm{M}_{\odot}}$, and the distributions from Sana et al. (2012); Oh et al. (2015) for stars with $M>5\mathrel{\mathrm{M}_{\odot}}$. For simulations with $\mathrm{f}_{\mathrm{bin}}=10\%$, we employed a log-uniform distribution for $M<5\mathrel{\mathrm{M}_{\odot}}$ and the Sana et al. (2012) period distribution for $M>5\mathrel{\mathrm{M}_{\odot}}$. Additionally, for the low binary fraction simulations, we disabled eigenevolution (Belloni et al. 2017). The separation was limited to the maximal value of $100\mathrel{\mathrm{AU}}$
In a part of our simulations (marked with NF in label), we introduced several modifications to investigate the influence of multiple stellar populations (see Giersz et al. 2024, for details). Specifically, we implemented a delay time ($t_{\rm delay}$) of $100\mathrel{\mathrm{Myr}}$ for the second population, during which these stars act as gas particles without undergoing stellar evolution, relaxation, or dynamical interactions. The gas accumulation for the second population begins at $t_{\rm delay\_fraction}=0.5$, meaning halfway through the delay period, coinciding with the asymptotic giant branch (AGB) wind contribution phase. To accommodate this, we activated AGB wind accumulation for the second population formation. The cluster’s orbital evolution was modified by implementing an artificial orbit change at $t=1000\mathrel{\mathrm{Myr}}$, where the galactocentric distance increases by a factor of 2 (see Giersz et al. 2024, for details). Finally, we employed single-population scaling, where the scaling parameters are applied exclusively to the first population, while the second population’s velocities are independently normalized. First and second populations are separately in the virial equilibrium.
Except dynamical evolution parameters, our simulations incorporate several key prescriptions for stars and binaries, some of them particularly relevant to ULX formation, specifically: We employed the rapid supernova mechanism as described by Fryer et al. (2012) for both neutron star and black hole formation; Neutron star and black hole natal kicks were drawn from a Maxwellian distribution with $\sigma=265\mathrel{\mathrm{km}/\mathrm{s}}$ (Hobbs et al. 2005); For black holes, these kicks were scaled by the fallback factor, which depends on the supernova prescription; Neutron stars formed through electron-capture supernovae were assigned substantially lower kicks with $\sigma=3\mathrel{\mathrm{km}/\mathrm{s}}$; Common envelope evolution was modeled using the BSE parameters $\alpha=0.5$ and $\lambda=0.0$. All simulated models in this study were set to a fixed metallicity of $\mathrm{Z}=0.001$.
More details on the MOCCA parameters can be found in Kamlah et al. (2022); Hypki et al. (2022, 2025); Giersz et al. (2024).
3 Results
3.1 Number evolution of ULXs
Figure 2: Evolution of ULX numbers in globular clusters. Expected number of ULXs ($\mathbb{E}[\mathcal{N}_{\mathrm{ULX}}]$) as a function of time since GC formation. All simulations with non-zero expected rates are presented. The line indicates the number of simulations with non-zero predictions (i.e. number of dots in this time bin). Each time bin spans $300\;\mathrel{\mathrm{Myr}}$ and the points are centered on these bins.
Figure 2 illustrates the temporal evolution of the expected number of ULXs [^3] Defined as the number of ULXs weighted by their observational probability in specific evolutionary time ranges. For example, an expected value of $0.001$ implies that, on average, one active ULX is observed across $1000$ simulations. Conversely, an expected value of $10$ indicates that, on average, $10$ ULXs are observed in a single simulation. This metric provides a probabilistic interpretation of ULX occurrence across simulations, accounting for their likelihood within evolutionary contexts. ($\mathbb{E}[\mathcal{N}_{\mathrm{ULX}}]$) across all simulations, accompanied by the count of simulations yielding non-zero $\mathbb{E}[\mathcal{N}_{\mathrm{ULX}}]$ for each time bin. The highest $\mathbb{E}[\mathcal{N}_{\mathrm{ULX}}]$ is observed in the initial evolutionary stages, typically within the first $300\mathrel{\mathrm{Myr}}$ (first bin in Figure 2). This trend persists across all simulations, irrespective of parameter variations. These findings align with previous studies on field populations, where dynamical interactions are negligible, which reported that the ULX phase predominantly occurs in the early evolutionary stages of binaries (e.g., Wiktorowicz et al. 2017). Observational evidence also supports this conclusion (e.g. Wolter et al. 2018).
This result suggests either:
  • Early ULX evolution is largely independent of cluster dynamics parameters, or
  • Insufficient time has elapsed for dynamics to significantly influence ULX formation.
The number of simulations with $\mathbb{E}[\mathcal{N}_{\mathrm{ULX}}]\neq 0$ decreases over time, resembling predictions for populations formed in burst-like star formation episodes (e.g. Wiktorowicz et al. 2017). In GCs, this trend may result from a combination of system age effects on ULX formation (as observed in field populations) and the interplay of positive and negative effects related to dynamics. Generally, our results indicate a higher likelihood of observing ULXs in younger stellar clusters compared to older ones.
[…]
The variations in the otherwise monotonically decreasing number of non-zero predictions can be attributed to randomness. The peak at $5\mathrel{\mathrm{Gyr}}$ and even more pronounced peak at $12\mathrel{\mathrm{Gyr}}$ are not correlated to any significant changes in the GC structure and are not statistically significant, therefore can be treated as anomalies.
Please note that the provided $\mathbb{E}[\mathcal{N}_{\mathrm{ULX}}]$ values represent predictions per cluster. To derive observational expectations, one must sum these values across all observed clusters and account for observational limitations, such as magnitude limits and resolution constraints. A detailed discussion of observational predictions will be addressed in a separate study.
3.2 Temporal Evolution and Parameter Dependence of $\mathbb{E}[\mathcal{N}_{\mathrm{ULX}}]$
Figure 3: Comprehensive overview of the $\mathbb{E}[\mathcal{N}_{\mathrm{ULX}}]$ evolution since the formation of a globular cluster (GC). Each panel represents a set of simulations with identical parameters: number of populations ($\mathrm{N}_{\mathrm{pop}}$), concentration ($\mathrm{R}_{\mathrm{h,2}}/\mathrm{R}_{\mathrm{h,1}}$), binary fraction ($\mathrm{f}_{\mathrm{bin}}$), and the presence of new features (NF). Within each panel, ULX populations from various environments are presented: tidally filling (TF) GC, non-tidally filling (nTF) GC, nodyn (binaries evolving without dynamical interactions), and escapers (progenitors formed in a GC, but experiencing the ULX phase after ejection from the cluster). See Table 1 for simulations parameters overview. Bars and points represent the $\mathbb{E}[\mathcal{N}_{\mathrm{ULX}}]$ in $300\mathrel{\mathrm{Myr}}$ bins
Figure 3 presents a comprehensive analysis of $\mathbb{E}[\mathcal{N}_{\mathrm{ULX}}]$ dependence on model parameters and its temporal evolution. The most significant effects are observed in parameters related to the environment, specifically whether the GC is tidally filling (TF) or non-tidally filling (nTF), and the general impact of dynamics (in-cluster vs. field formation). These results support the hypothesis that dynamical interactions strongly influence the formation of ULXs in dense stellar environments. In particular, the exchange process may lead to the formation of binaries that are not possible through regular formation channels. The specific trends and patterns are described below.
Non-tidally filling clusters exhibit significantly larger ULX populations compared to tidally filling clusters. This discrepancy can be attributed to the extended time frame density available for dynamical interactions to enhance progenitor formation in non-tidally filling clusters, whereas tidally filling clusters experience more early ejections. More generally, the expected number of ULXs formed in GCs is substantially higher than in corresponding field populations, particularly for non-tidally filling clusters. Typically, ULXs formed in GCs demonstrate higher $\mathbb{E}[\mathcal{N}_{\mathrm{ULX}}]$ rates in specific time bins, and these rates are more continuous, whereas field populations mostly exhibit isolated occurrences throughout the timeline. These relations do not apply to early evolutionary stages where all simulations show high and similar values of $\mathbb{E}[\mathcal{N}_{\mathrm{ULX}}]$ (see Section 3.1).
Simulations with two populations (middle column) generally produce more ULXs than those with a single population (left column). This can be attributed to the higher concentration of the second population, which enhances ULX progenitor formation. However, when the concentration is excessively high (right column), this effect is less pronounced, which supports the claim that too frequent and strong dynamical interactions can substantially destroy ULX progenitors.
The initial binary fraction appears to have a limited influence on ULX formation in GCs. For $\mathrm{f}_{\mathrm{bin}}=95\%$ wide binaries are quickly destroyed in dynamical interactions and mostly remain binaries similar to ones for $\mathrm{f}_{\mathrm{bin}}=10\%$. The population of ULX progenitors remains sparse and emerges regardless of the initial relative number of binaries. Dynamical effects effectively destroy existing binaries and create new ones, a process primarily responsible for the formation of ULX progenitors at later evolutionary times. Conversely, the binary fraction influences field populations, with a positive correlation between initial binaries and ULX progenitors, as there are no external processes to destroy ULX progenitors or create new ones in isolation.
The incorporation of new features (NF) has a noticeable effect on two-population simulations with moderate concentration ($\mathrm{R}_{\mathrm{h,2}}/\mathrm{R}_{\mathrm{h,1}}=0.2$; Figure 3, middle column). Simulations with NF exhibit higher $\mathbb{E}[\mathcal{N}_{\mathrm{ULX}}]$ values compared to those without NF (nNF). The NF do not affect field populations as they primarily relate to the time delay for the second population.
The escaper populations generally align with the ULX population formed in situ, suggesting ongoing ejection (relacation and interactions), and thus a continuous supply of ULX progenitors to the escaper population, which later become ULXs while unbound to the cluster. There are generally fewer ULXs among escapers than in in-cluster populations. However, ULXs from escapers can dominate field populations. In general, there are no significant differences in the formation time of the ULXs from escapers to those in the field. An exception occurs in simulations with relatively strong ULX formation ($\mathrm{N}_{\mathrm{pop}}=2$, $\mathrm{R}_{\mathrm{h,2}}/\mathrm{R}_{\mathrm{h,1}}=0.2$), where ULXs from escapers form at evolutionary ages not covered by field populations. For tidally filling clusters, the rate of ULX formation from escapers can exceed that of in-cluster formation, especially in simulations without new features (nNF).
3.3 Formation of ULXs
This section focuses on ULX progenitors - systems that will evolve into ULXs - and their evolution up to the initial ULX phase. It is important to note that some systems may experience multiple ULX phases interspersed with periods of quiescence or XRB phases ($\mathrm{L}_{\mathrm{X}}<10^{39}\mathrel{\mathrm{erg}/\mathrm{s}}$). Additionally, individual stars can participate in multiple ULXs, each with a different companion. This phenomenon is particularly prevalent for IMBHs (see Sect. 3.4.1).
Table 2: Progenitor properties: ZAMS progenitors, ULX onset time, and total ULXs for different configurations.
First Population Second Population
binary single binary single
$\mathrm{N}_{\mathrm{prog}}$
A-D(p) $148.67\pm 92.34\;(36)$ $5.86\pm 5.75\;(14)$
A-n $9.29\pm 15.49\;(14)$ $28.75\pm 17.07\;(24)$ $8.00\pm 13.75\;(9)$ $5.55\pm 8.08\;(11)$
D-A(p) $2.22\pm 2.28\;(9)$
D-n $23.31\pm 24.33\;(13)$ $27.96\pm 18.79\;(24)$ $12.33\pm 17.78\;(15)$ $8.18\pm 8.44\;(11)$
n-A $6.73\pm 9.08\;(11)$ $4.57\pm 6.11\;(7)$
n-D $12.25\pm 7.31\;(12)$ $6.54\pm 4.82\;(13)$
$\mathrm{med}(\mathrm{t}_{\mathrm{ULX,start}})$
A-D(p) $6.21\pm 0.44\;(36)$ $297.78\pm 209.46\;(14)$
A-n $3275.22\pm 3442.16\;(14)$ $345.72\pm 1209.67\;(24)$ $1726.87\pm 3277.15\;(9)$ $1922.14\pm 3139.30\;(11)$
D-A(p) $6924.70\pm 5695.35\;(9)$
D-n $3752.29\pm 2056.81\;(13)$ $334.86\pm 1242.41\;(24)$ $2074.99\pm 1919.42\;(15)$ $1599.31\pm 2702.66\;(11)$
n-A $3291.82\pm 3514.45\;(11)$ $1851.74\pm 2567.34\;(7)$
n-D $3626.39\pm 2618.49\;(12)$ $1596.63\pm 2707.93\;(13)$
$\mathbb{E}[\mathcal{N}_{\mathrm{ULX}}]_{total}$
A-D(p) $88.39\pm 162.02\;(36)$ $0.32\pm 0.37\;(14)$
A-n $56.30\pm 82.00\;(14)$ $77.75\pm 147.37\;(24)$ $55.16\pm 82.91\;(9)$ $7.79\pm 15.10\;(11)$
D-A(p) $42.20\pm 123.25\;(9)$
D-n $68.97\pm 132.00\;(13)$ $49.56\pm 94.57\;(24)$ $22.28\pm 28.24\;(15)$ $26.22\pm 62.54\;(11)$
n-A $49.22\pm 68.55\;(11)$ $2.06\pm 1.91\;(7)$
n-D $91.36\pm 174.29\;(12)$ $3.01\pm 2.27\;(13)$
444 Number of progenitors on zams ($\mathrm{N}_{\mathrm{prog}}$), median fo the time when the ULX phase commences ($\mathrm{med}(\mathrm{t}_{\mathrm{ULX,start}})$), and the total (i.e. through the entire history) expected number of ULXs ($\mathbb{E}[\mathcal{N}_{\mathrm{ULX}}]_{total}$). Number in parenthesis inform about the number of simulations with non-zero $\mathbb{E}[\mathcal{N}_{\mathrm{ULX}}]$ of specific configuration. Results are present for different configurations of progenitors: A stands for the progenitor of an accretor, whereas D for the progenitor of donor, n stands for stars which do not become part of a ULX, or is a placeholder in case of single stars. p stands for pristine and means that there were no exchanges in a binary before ULX formation. First symbol represents the primary (i.e. more massive star on ZAMS), whereas the second letter the secondary. Results are divided into First and Second Populations and also between binaries and single stars. Values represent the mean and one sigma error for simulations in which the specific ULX where present.
Table 2 provides a comprehensive overview of ULX progenitors on ZAMS, i.e. stars that later became components of ULXs. The most common progenitors are those of pristine ULXs, i.e. binaries that do not undergo exchanges or disruptions prior to the ULX phase. In these systems, the primary typically becomes the accretor and the secondary the donor (A-D(p)). These progenitors evolve rapidly, initiating the ULX phase after approximately $\mathrm{med}(\mathrm{t}_{\mathrm{ULX,start}})\approx 6\mathrel{\mathrm{Myr}}$, in stark contrast to other progenitors which generally begin after $\mathrm{med}(\mathrm{t}_{\mathrm{ULX,start}})\gtrsim 3\mathrel{\mathrm{Gyr}}$. Pristine ULXs are observed in all simulations.
Interestingly, some pristine ULXs form where the secondary, less massive on ZAMS, becomes the accretor (D-A(p)). These progenitors are considerably rarer (about two orders of magnitude less common than A-D(p)) and typically form ULXs only in later evolutionary phases ($\sim 7\mathrel{\mathrm{Gyr}}$).
Non-pristine ULXs can originate from both binary and single star progenitors on ZAMS. However, stars born in binaries more frequently become components of ULXs, even in non-pristine cases. These systems invariably undergo significant interactions, such as exchanges or disruptions, before forming a ULX. While some of these ULXs can form relatively early ($\sim 100\mathrel{\mathrm{Myr}}$), the majority emerge in later evolutionary phases ($>3\mathrel{\mathrm{Gyr}}$). In non-pristine ULXs, both the primary and secondary can act as either the accretor or donor with comparable probability. Our simulations did not reveal any non-pristine ULXs where both the primary and secondary are parts of ULXs (i.e., ”A-D” or ”D-A” type), though this is likely due to limited statistics rather than a physical constraint.
3.4 ULX properties
In this section we discuss the properties of in-cluster ULXs, i.e. these which reside inside the cluster. For discussion of ULXs among escapers see Section 3.7.
Table 3: Properties of ULXs by Group
group $\mathrm{k}_{\mathrm{acc}}$ $\mathrm{k}_{\mathrm{don}}$ $\mathrm{M}_{\mathrm{acc}}$ $\mathrm{M}_{\mathrm{don}}$ $a$ $e$ $\log_{10}\mathrm{L}_{\mathrm{X,max}}$ $\Delta\mathrm{t}_{\mathrm{ULX}}$ $\mathrm{t}_{\mathrm{phys,min}}$
$G_{\textrm{BH,MS}}$ BH MS $27.64^{+18.02}_{-19.43}$ $45.72^{+81.05}_{-42.83}$ $40.05^{+104.24}_{-32.82}$ $0.00^{+0.58}_{-0.00}$ $40.36^{+0.38}_{-1.11}$ $0.41^{+4.71}_{-0.41}$ $6.16^{+380.87}_{-2.86}$
$G_{\textrm{BH,HG}}$ BH HG $22.70^{+22.73}_{-14.22}$ $56.64^{+46.70}_{-51.47}$ $119.21^{+905.18}_{-88.52}$ $0.00^{+0.45}_{-0.00}$ $40.35^{+0.32}_{-0.84}$ $0.13^{+1.35}_{-0.13}$ $5.63^{+131.94}_{-2.74}$
$G_{\textrm{BH,CHeB}}$ BH CHeB $16.88^{+29.46}_{-8.38}$ $32.52^{+43.80}_{-18.12}$ $233.71^{+2887.84}_{-179.57}$ $0.03^{+0.36}_{-0.03}$ $40.18^{+0.38}_{-0.60}$ $0.10^{+0.34}_{-0.09}$ $7.52^{+9.56}_{-4.29}$
$G_{\textrm{IMBH}}$ BH * $9351.31^{+4162.72}_{-8738.78}$ $0.96^{+168.06}_{-0.73}$ $28.54^{+3054.93}_{-28.26}$ $0.56^{+0.39}_{-0.56}$ $40.84^{+2.29}_{-1.76}$ $0.00^{+3.06}_{-0.00}$ $2207.25^{+11595.73}_{-2195.15}$
$G_{\textrm{NS}}$ NS ** $1.27^{+1.17}_{-0.16}$ $0.78^{+4.51}_{-0.51}$ $0.45^{+51.06}_{-0.43}$ $0.00^{+0.34}_{-0.00}$ $39.22^{+0.52}_{-0.21}$ $0.05^{+27.48}_{-0.05}$ $331.03^{+12959.06}_{-258.73}$
other NS or BH *** $22.13^{+22.42}_{-20.84}$ $3.22^{+32.80}_{-2.89}$ $489.01^{+3081.72}_{-488.99}$ $0.00^{+0.64}_{-0.00}$ $39.95^{+0.87}_{-0.85}$ $0.06^{+32.44}_{-0.06}$ $143.70^{+13216.21}_{-138.87}$
555 Median properties and $95\%$ confidence intervals for ULX groups identified in this study. Column definitions: $\mathrm{k}_{\mathrm{acc}}$ accretor type; $\mathrm{k}_{\mathrm{don}}$ donor type; $\mathrm{M}_{\mathrm{acc}}$ accretor mass ($\mathrel{\mathrm{M}_{\odot}}$); $\mathrm{M}_{\mathrm{don}}$ donor mass ($\mathrel{\mathrm{M}_{\odot}}$); $a$ orbital separation ($\mathrel{\mathrm{R}_{\odot}}$); $e$ eccentricity; $\mathrm{L}_{\mathrm{X,max}}$ maximum X-ray luminosity ($\mathrel{\mathrm{erg}/\mathrm{s}}$); $\Delta\mathrm{t}_{\mathrm{ULX}}$ ULX lifetime ($\mathrel{\mathrm{Myr}}$); $\mathrm{t}_{\mathrm{phys,min}}$ ULX phase onset age. Donor types: MS - Main Sequence, HG - Hertzsprung Gap, RG - Red Giant, CHeB - Core Helium Burning. For groups with mixed donor types: * HG: $47\%$, MS: $41\%$, RG: $6\%$, CHeB: $3\%$; ** RG: $83\%$, HG: $7\%$, MS: $6\%$; *** RG: $75\%$, CHeB: $21\%$; other companion types represent less then $5\%$ of all companions.
[…]
NS accretors in ULXs ($G_{\textrm{NS}}$) have median masses around $1.27\mathrel{\mathrm{M}_{\odot}}$. These systems typically have lower X-ray luminosities ($\log_{10}\mathrm{L}_{\mathrm{X,max}}\approx 39.22$) compared to their BH counterparts, but still clearly exceed the Eddington limit for a typical NS.
The orbital characteristics of ULXs vary significantly across different groups. $G_{\textrm{BH,MS}}$ systems have median separations of $40.05\mathrel{\mathrm{R}_{\odot}}$, while $G_{\textrm{NS}}$ systems show much tighter orbits (median $0.45\mathrel{\mathrm{R}_{\odot}}$). Most ULXs in our sample have low eccentricities, suggesting that tidal forces or a common envelope phase have circularized their orbits before the onset of the ULX phase.
Duty cycles vary widely among ULX groups, with $G_{\textrm{BH,MS}}$ systems showing the highest median value ($0.41$), while other groups have much lower values. This variability in duty cycles may explain the transient nature observed in many ULXs and has implications for their detectability (cf. Wiktorowicz et al. 2017).
A notable subset of our sample consists of IMBH-powered systems ($G_{\textrm{IMBH}}$). These ULXs are characterized by extremely high accretor masses (median $9351\mathrel{\mathrm{M}_{\odot}}$) and the highest X-ray luminosities in our sample ($\log_{10}\mathrm{L}_{\mathrm{X,max}}\approx 40.84$). Such luminous systems, called extreme-ULXs (Wiktorowicz et al. 2019), are one of the main observational candidates for IMBHs and can provide insights into their formation and properties.
The diversity in donor types across ULX groups highlights the various evolutionary pathways that can lead to ULX formation. MS, HG, and RG donors are all well-represented, suggesting that ULXs can form and persist across a wide range of stellar evolutionary stages
Table 4: Fraction of ULXs in Different Groups Across Simulations
label $G_{\textrm{BH,MS}}$ $G_{\textrm{BH,HG}}$ $G_{\textrm{BH,CHeB}}$ $G_{\textrm{IMBH}}$ $G_{\textrm{NS}}$ other
npop1-fb10-TF $0.17$ $0.33$ $0.36$ $0.11$ $0.04$
npop1-fb10-nTF $0.16$ $0.32$ $0.30$ $0.12$ $0.11$
npop1-fb10-nTF-nodyn $0.18$ $0.36$ $0.37$ $0.05$ $0.04$
npop1-fb95-TF $0.11$ $0.40$ $0.33$ $0.12$ $0.04$
npop1-fb95-nTF $0.12$ $0.39$ $0.28$ $0.11$ $0.11$
npop1-fb95-nTF-nodyn $0.11$ $0.43$ $0.31$ $0.11$ $0.04$
npop2-cpop05-fb10-TF $0.09$ $0.39$ $0.29$ $0.09$ $0.13$
npop2-cpop05-fb10-nTF $0.07$ $0.29$ $0.20$ $0.33$ $0.04$ $0.06$
npop2-cpop05-fb10-nTF-nodyn $0.10$ $0.46$ $0.32$ $0.04$ $0.07$
npop2-cpop05-fb10-TF-NF $0.15$ $0.38$ $0.27$ $0.05$ $0.14$
npop2-cpop05-fb10-nTF-NF $0.06$ $0.25$ $0.16$ $0.44$ $0.05$ $0.03$
npop2-cpop05-fb10-nTF-nodyn-NF $0.11$ $0.48$ $0.33$ $0.08$
npop2-cpop05-fb95-TF $0.11$ $0.29$ $0.42$ $0.11$ $0.06$
npop2-cpop05-fb95-nTF $0.11$ $0.25$ $0.35$ $0.16$ $0.07$ $0.06$
npop2-cpop05-fb95-nTF-nodyn $0.12$ $0.29$ $0.41$ $0.14$ $0.04$
npop2-cpop05-fb95-TF-NF $0.13$ $0.27$ $0.44$ $0.10$ $0.06$
npop2-cpop05-fb95-nTF-NF $0.10$ $0.18$ $0.29$ $0.34$ $0.05$ $0.04$
npop2-cpop05-fb95-nTF-nodyn-NF $0.12$ $0.30$ $0.43$ $0.12$ $0.04$
npop2-cpop2-fb10-TF $0.10$ $0.41$ $0.30$ $0.09$ $0.10$
npop2-cpop2-fb10-nTF $0.06$ $0.28$ $0.19$ $0.36$ $0.04$ $0.07$
npop2-cpop2-fb10-nTF-nodyn $0.10$ $0.46$ $0.32$ $0.04$ $0.07$
npop2-cpop2-fb10-TF-NF $0.12$ $0.39$ $0.30$ $0.09$ $0.09$
npop2-cpop2-fb10-nTF-NF $0.16$ $0.16$ $0.09$ $0.41$ $0.18$
npop2-cpop2-fb10-nTF-nodyn-NF $0.11$ $0.48$ $0.33$ $0.08$
npop2-cpop2-fb95-TF $0.13$ $0.26$ $0.44$ $0.13$ $0.05$
npop2-cpop2-fb95-nTF $0.10$ $0.24$ $0.31$ $0.20$ $0.09$ $0.06$
npop2-cpop2-fb95-nTF-nodyn $0.12$ $0.29$ $0.42$ $0.13$ $0.04$
npop2-cpop2-fb95-TF-NF $0.12$ $0.28$ $0.45$ $0.11$ $0.05$
npop2-cpop2-fb95-nTF-NF $0.13$ $0.17$ $0.22$ $0.34$ $0.14$
npop2-cpop2-fb95-nTF-nodyn-NF $0.12$ $0.29$ $0.41$ $0.14$ $0.04$
666 Table presents the fractional distribution of ULXs across different groups for various simulations. For group definitions see Table 3
Huge majority are ULXs with BH accretors (typically $>80\%$) formed early in cluster evolution.
Table 4 shows a modest anticorrelation between the relative size of the $G_{\textrm{IMBH}}$ and BH ULXs (Kendall [^7] Kendall’s $\tau$ is preferable to Pearson’s $r$ when working with small samples and non-normal distributions, while still providing a similarly interpretable measure of association between $-1$ and $+1$.’s $\tau=-0.49,-0.38,-0.37$, for $G_{\textrm{BH,MS}}$, $G_{\textrm{BH,HG}}$, and $G_{\textrm{BH,CHeB}}$, respectively). The presence of IMBH results in a continuous formation of IMBH powered ULX though the cluster history which, being an independent sourse to other BH ULX groups present mostly in the early cluster evolution phases, lowers their fractional imput to the ULX population. IMBH quickly removes BHs and possibly NSs from the system (e.g. Hong et al. 2020).
3.4.1 IMBH ULXs
In this study, we define intermediate-mass black holes (IMBHs) as those with masses exceeding $500\mathrel{\mathrm{M}_{\odot}}$. IMBH ULXs were observed in 6 simulations, all featuring multiple stellar populations (npop2) and non-tidally filling initial conditions (nTF). When present, IMBHs contribute significantly to the ULX population, accounting for 16-44% of ULXs over a Hubble time (Table 4).
IMBHs form rapidly, within several Myr after the zero-age main sequence (ZAMS), classifying them as ”fast IMBHs” (Giersz et al. 2015). In simulations with delayed second populations, IMBH formation occurs 100+ Myr after the initial population. Notably, IMBH ULXs form exclusively in non-tidally filling (nTF) simulations, strenghtening the preference for denser stellar environments.
Fisher’s exact tests revealed no significant correlations between simulation parameters and IMBH formation at the 90% confidence level. However, when considering only simulations with multiple populations (pop=2) and active dynamics, a statistically significant relationship (95% confidence) emerged between the filling parameter and IMBH ULX presence, favoring non-tidally filling clusters.
Table 5: IMBH ULX properties
label $\mathrm{t}_{\mathrm{phys,min}}$ $\mathrm{t}_{\mathrm{phys,max}}$ duty-cycle [%] companions
npop2-cpop05-fb10-nTF $7.95$ $13,306.24$ $0.06$ MS: 22, HeWD: 4, COWD: 3, TPAGB: 2, CHeB: 2
npop2-cpop05-fb10-nTF-NF $100.63$ $14,925.72$ $0.02$ MS: 41, COWD: 5, RG: 3, HeWD: 2, CHeB: 1,
EAGB: 1, TPAGB: 1, ONeWD: 1
npop2-cpop05-fb95-nTF $11.49$ $13,791.13$ $0.87$ MS: 29, COWD: 13, HG: 1, HeWD: 1, RG: 1
npop2-cpop05-fb95-nTF-NF $104.89$ $14,751.19$ $0.22$ MS: 89, COWD: 12, TPAGB: 2, HeWD: 2, CHeB: 2,
HeGB: 1, ONeWD: 1, HG: 1
npop2-cpop2-fb10-nTF $9.19$ $14,171.06$ $0.14$ MS: 32, COWD: 5, HeWD: 2, RG: 1, TPAGB: 1,
ONeWD: 1
npop2-cpop2-fb95-nTF $8.99$ $14,546.35$ $0.43$ MS: 43, COWD: 11, HG: 3, HeWD: 3, TPAGB: 2,
RG: 2
888 Table presents the properties of $6$ IMBH which appeared in our simulations.
[…]
IMBH ULXs formed exclusively in simulations with two stellar populations and non-tidally filling initial conditions. The centrally concentrated second population facilitates stellar interactions and mergers, while nTF clusters retain more stars, particularly black holes, in central regions, extending the IMBH formation window.
3.5 ULX progenitors
The initial properties of binary systems that evolve into ULXs provide crucial insights into their formation channels and evolutionary pathways. Our results indicate that the primary distinguishing factor between ULX progenitors and the general stellar population is their higher initial mass. No other significant constraints on ZAMS properties were identified for ULX progenitors.
Table 6: Initial Properties of ULX Progenitors by Group
group * $\mathrm{M}_{\mathrm{ZAMS}}$ $\mathrm{f}_{\mathrm{esc}}$ $\mathrm{f}_{\mathrm{bin}}$ mode $\mathrm{f}_{\mathrm{prim}}$ $\mathrm{f}_{\mathrm{pris}}$ $\mathrm{M}_{\mathrm{comp}}$ $a$ $e$
$G_{\textrm{BH,MS}}$ A $47.00^{+73.39}_{-46.10}$ $0.07$ $0.91$ A-D(p) (85%) $0.98$ $0.94$ $20.89^{+64.91}_{-12.95}$ $71.01^{+3187.79}_{-39.51}$ $0.19^{+0.67}_{-0.19}$
D $18.48^{+59.00}_{-17.59}$ $0.07$ $0.91$ A-D(p) (83%) $0.07$ $0.91$ $48.97^{+71.42}_{-47.11}$ $67.00^{+3191.81}_{-42.35}$ $0.18^{+0.68}_{-0.18}$
$G_{\textrm{BH,HG}}$ A $48.13^{+97.30}_{-40.02}$ $0.03$ $0.96$ A-D(p) (95%) $0.99$ $0.98$ $33.22^{+78.41}_{-24.70}$ $115.21^{+2330.09}_{-79.31}$ $0.13^{+0.60}_{-0.13}$
D $32.29^{+79.34}_{-30.81}$ $0.03$ $0.97$ A-D(p) (94%) $0.01$ $0.98$ $48.59^{+96.84}_{-27.64}$ $114.91^{+2149.83}_{-79.93}$ $0.14^{+0.59}_{-0.14}$
$G_{\textrm{BH,CHeB}}$ A $37.02^{+76.30}_{-36.45}$ $0.11$ $0.89$ A-D(p) (88%) $1.00$ $0.99$ $23.69^{+48.14}_{-12.46}$ $398.98^{+3919.72}_{-350.69}$ $0.21^{+0.60}_{-0.21}$
D $21.93^{+49.90}_{-21.37}$ $0.11$ $0.89$ A-D(p) (88%) $0.01$ $0.99$ $37.98^{+75.81}_{-16.70}$ $398.98^{+3897.78}_{-352.62}$ $0.21^{+0.60}_{-0.21}$
$G_{\textrm{IMBH}}$ A $13.54^{+2.57}_{-3.50}$ $0.00$ $0.67$ A-n (100%) $1.00$ $0.00$ $9.67^{+2.20}_{-0.85}$ $75.45^{+1.96}_{-39.57}$ $0.06^{+0.07}_{-0.05}$
D $1.19^{+14.35}_{-1.05}$ $0.00$ $0.88$ D-n (61%) $0.56$ $0.00$ $0.90^{+11.26}_{-0.72}$ $36.70^{+276433.30}_{-30.94}$ $0.34^{+0.58}_{-0.33}$
$G_{\textrm{NS}}$ A $5.98^{+12.30}_{-5.87}$ $0.39$ $0.68$ A-D(p) (49%) $0.90$ $0.74$ $3.04^{+8.86}_{-2.94}$ $1410.90^{+600401.60}_{-1393.38}$ $0.32^{+0.58}_{-0.32}$
D $2.07^{+8.32}_{-1.98}$ $0.39$ $0.70$ A-D(p) (48%) $0.27$ $0.71$ $5.46^{+11.13}_{-5.33}$ $1410.90^{+724659.10}_{-1400.20}$ $0.34^{+0.58}_{-0.34}$
other A $24.70^{+46.22}_{-24.56}$ $0.15$ $0.80$ A-D(p) (57%) $0.89$ $0.73$ $10.44^{+40.18}_{-10.24}$ $1849.40^{+226115.60}_{-1824.04}$ $0.22^{+0.70}_{-0.22}$
D $4.73^{+31.64}_{-4.61}$ $0.14$ $0.80$ A-D(p) (52%) $0.28$ $0.66$ $22.43^{+46.86}_{-22.31}$ $1409.40^{+338400.60}_{-1396.45}$ $0.31^{+0.59}_{-0.31}$
999 Values are provided as median with 95% confidence intervals or averages. Groups are defined in Table 3 and accompaning text. * denotes component (A - accretor, D - donor). $\mathrm{M}_{\mathrm{ZAMS}}$: ZAMS mass ($\mathrel{\mathrm{M}_{\odot}}$); esc: fraction of ULXs formed from escapers; bin: binary fraction among progenitors; mode: dominant formation mode (see Table 2); $\mathrm{f}_{\mathrm{prim}}$: fraction of progenitors formed from primary stars; $\mathrm{f}_{\mathrm{pris}}$: fraction of pristine ULXs formed; $\mathrm{M}_{\mathrm{comp}}$: companion ZAMS mass ($\mathrel{\mathrm{M}_{\odot}}$); $a$: ZAMS separation ($\mathrel{\mathrm{R}_{\odot}}$); $e$: ZAMS eccentricity. Mode, $\mathrm{f}_{\mathrm{prim}}$, $\mathrm{f}_{\mathrm{pris}}$, $\mathrm{M}_{\mathrm{comp}}$, $a$, and $e$ values are calculated only for binary progenitors.
Table 6 summarizes the ZAMS properties of ULX progenitors, categorized by ULX group and component role (accretor or donor). Accretor progenitors typically have higher ZAMS masses ($\mathrm{M}_{\mathrm{ZAMS}}$) than their companions ($\mathrm{M}_{\mathrm{comp}}$), consistent with being predominantly formed from primary stars ($\mathrm{f}_{\mathrm{prim}}\geq 0.89$).
The majority of ULX progenitors originate in binary systems ($\mathrm{f}_{\mathrm{bin}}\geq 0.67$), with most maintaining their original pairing to form the ULX ($\mathrm{f}_{\mathrm{pris}}\geq 0.66$, except for the $G_{\textrm{IMBH}}$ group). For accretors in the $G_{\textrm{IMBH}}$ and $G_{\textrm{NS}}$ groups, $\mathrm{f}_{\mathrm{bin}}\approx 0.66$, indicating no preference between single and binary star origins.
Initial separations ($a$) exhibit a highly skewed distribution, with $G_{\textrm{NS}}$ group progenitors typically having much wider separations than other groups. Values for the $G_{\textrm{IMBH}}$ group are not particularly informative as these stars do not become components of ULXs without undergoing strong dynamical interactions. All groups show a preference for lower initial eccentricities ($\mathrm{med}(e)\lesssim 0.3$).
The $G_{\textrm{IMBH}}$ group shows distinct characteristics compared to other groups. It has the lowest ZAMS masses for both accretors and donors, a unique dominant formation modes (A-n for accretors and D-n for donors), and the lowest $\mathrm{f}_{\mathrm{bin}}$. The $G_{\textrm{BH,MS}}$ and $G_{\textrm{BH,HG}}$ groups exhibit similar patterns, with high ZAMS masses and a strong preference for the A-D(p) formation mode. The $G_{\textrm{NS}}$ group has the highest fraction of ULXs formed from escapers ($\mathrm{f}_{\mathrm{esc}}=0.39$).
3.6 ULX descendants
Table 7: ULX descendants
D E X M O
npop1-fb10-TF $67$ $14$
npop1-fb10-nTF $4$ $22$ $46$ $25$ $4$
npop1-fb95-TF $222$ $36$ $1$
npop1-fb95-nTF $18$ $69$ $152$ $70$ $19$
npop2-cpop05-fb10-TF $2$ $35$ $11$ $13$
npop2-cpop05-fb10-TF-NF $2$ $25$ $21$ $14$ $2$
npop2-cpop05-fb10-nTF $18$ $20$ $10$ $42$ $1$
npop2-cpop05-fb10-nTF-NF $23$ $12$ $10$ $65$ $1$
npop2-cpop05-fb95-TF $4$ $117$ $30$ $37$ $3$
npop2-cpop05-fb95-TF-NF $13$ $90$ $49$ $37$ $9$
npop2-cpop05-fb95-nTF $55$ $72$ $38$ $81$ $3$
npop2-cpop05-fb95-nTF-NF $66$ $54$ $41$ $136$ $8$
npop2-cpop2-fb10-TF $37$ $2$ $12$ $3$
npop2-cpop2-fb10-TF-NF $3$ $24$ $15$ $14$ $1$
npop2-cpop2-fb10-nTF $17$ $14$ $16$ $55$ $2$
npop2-cpop2-fb10-nTF-NF $9$ $27$ $60$ $67$ $12$
npop2-cpop2-fb95-TF $2$ $115$ $17$ $34$ $9$
npop2-cpop2-fb95-TF-NF $4$ $88$ $42$ $36$ $13$
npop2-cpop2-fb95-nTF $73$ $41$ $105$ $66$
npop2-cpop2-fb95-nTF-NF $17$ $60$ $124$ $130$ $45$
101010 Distribution of ULX descendant fates for various simulation configurations. Abbreviations: D - disruption; E - escape; X - exchange; M - merger; O - other. Only simulations with dynamical interactions are included. The last row represents approximate averages across all simulations. See text for discussion and details.
The fate of ULX systems after their active phase provides valuable insights into their evolutionary pathways and potential contributions to other astrophysical phenomena. Table 7 summarizes the various outcomes for ULX systems across different simulation configurations.
ULX systems can undergo several fates: disruption, escape from the cluster, exchange interactions, mergers, or other outcomes. The distribution of these fates appears largely stochastic across different simulation parameters. Systems categorized as ”other” typically experience no significant events for the remainder of their evolution and are predominantly NS/BH + WD or DCOs at the end of simulation.
Merger events typically occur $0.79^{+0.74}_{-0.53}\mathrel{\mathrm{Myr}}$ after the ULX phase. Double compact objects (DCOs) form in $\sim 63\%$ of cases, with $\sim 34\%$ of them being ejected from the cluster before merger or the end of simulation. These DCOs that remain in the cluster may contribute to the population of gravitational wave (GW) sources with merger rate of $\sim 11\%$.
Disruptions may result from dynamical interactions or binary evolution. In exchanges, at least one of the components continues evolution as part of some other binary. Our results suggest, that in initially more centrally concentrated environment (nTF) there are more disruptions and exchanges after the ULX phases, then in other simulations.
3.7 ULXs among escapers
Cluster-formed binaries are frequently ejected, suggesting that some field ULXs may originate from GCs. These ULX progenitors become unbound from their parent clusters and initiate the ULX phase while residing in the galactic field, evolving as isolated binaries.
Figure 3 illustrates the expected numbers and ages of ULXs among escapers. Generally, escaper ULXs coexist in similar numbers and at comparable evolutionary times as their in-cluster counterparts, except the very early phase of cluster evolution ($\lesssim 300\mathrel{\mathrm{Myr}}$). Table 8 presents a detailed comparison of properties between in-cluster and escaper ULXs.
Table 8: Properties of Escaper and In-cluster ULXs
escaper in-cluster
NS BH NS BH
count $12.50^{+25.12}_{-10.55}$ $19.00^{+20.35}_{-15.05}$ $7.00^{+89.20}_{-6.00}$ $142.50^{+158.18}_{-88.50}$
Typ. companion HeMS (70%) CHeB (80%) HG (74%) HG (55%)
$\mathrm{med}(\mathrm{M}_{\mathrm{acc}})$ $1.32^{+0.52}_{-0.18}$ $16.58^{+16.05}_{-1.98}$ $1.26^{+0.34}_{-0.09}$ $22.51^{+147.80}_{-4.10}$
$\mathrm{med}(\mathrm{M}_{\mathrm{don}})$ $0.54^{+1.87}_{-0.06}$ $19.00^{+15.82}_{-15.04}$ $1.80^{+6.95}_{-1.41}$ $40.55^{+7.59}_{-23.52}$
$\mathrm{med}(a)$ $0.43^{+14.08}_{-0.40}$ $126.33^{+89.54}_{-88.43}$ $16.55^{+18.52}_{-16.51}$ $123.62^{+76.99}_{-76.17}$
$\mathrm{med}(e)$ $0.00^{+0.11}_{-0.00}$ $0.09^{+0.26}_{-0.09}$ $0.01^{+0.21}_{-0.01}$ $0.00^{+0.03}_{-0.00}$
$\mathrm{med}(\log_{10}\mathrm{L}_{\mathrm{X,max}})$ $39.22^{+0.24}_{-0.10}$ $40.05^{+0.09}_{-0.06}$ $39.20^{+0.17}_{-0.19}$ $40.28^{+0.11}_{-0.14}$
$\mathrm{med}(\Delta\mathrm{t}_{\mathrm{ULX}})$ $0.12^{+0.08}_{-0.11}$ $0.12^{+0.18}_{-0.08}$ $0.05^{+10.55}_{-0.05}$ $0.11^{+0.02}_{-0.07}$
$t_{\mathrm{ULX,min}}$ $115.30^{+120.96}_{-73.08}$ $3.49^{+4.99}_{-0.60}$ $119.45^{+1512.99}_{-106.11}$ $2.85^{+0.48}_{-0.13}$
$t_{\mathrm{ULX,max}}$ $5767.74^{+9100.42}_{-5558.67}$ $1994.71^{+11988.23}_{-1953.63}$ $11682.46^{+3038.02}_{-11616.14}$ $11507.08^{+3391.92}_{-11440.52}$
111111 Comparison of properties between ULXs in-cluster and those among escapers. Values represent the median with errors indicating the 10th and 90th percentiles of the all simulations with dynamics included. Most of the values represent the median values for the simulation ($\mathrm{med}()$). $\mathrm{M}_{\mathrm{acc}}$/$\mathrm{M}_{\mathrm{don}}$ accretor/donor mass [$\mathrel{\mathrm{M}_{\odot}}$], $a$ orbital separation [$\mathrel{\mathrm{R}_{\odot}}$], $e$ eccentricity, $\mathrm{L}_{\mathrm{X,max}}$ maximum X-ray luminosity [$\mathrel{\mathrm{erg}/\mathrm{s}}$], $t_{\mathrm{ULX,min}}$/$t_{\mathrm{ULX,max}}$ minimum/maximum age of ULX phase onset [$\mathrel{\mathrm{Myr}}$].
Escaper ULXs exhibit several distinct characteristics compared to their in-cluster counterparts. They tend to have lower-mass accretors and donors, with median masses of $13.00$ and $6.05\mathrel{\mathrm{M}_{\odot}}$, respectively, compared to $22.24$ and $39.34\mathrel{\mathrm{M}_{\odot}}$ for in-cluster ULXs. Escaper ULXs also have tighter orbits (median $57.42$ vs. $119.93\mathrel{\mathrm{R}_{\odot}}$) and a significantly higher fraction of neutron star accretors ($40\%$ vs. $4\%$). The median maximum X-ray luminosity of escaper ULXs ($\log_{10}\mathrm{L}_{\mathrm{X,max}}=39.89\mathrel{\mathrm{erg}/\mathrm{s}}$) is slightly lower than that of in-cluster ULXs ($\log_{10}\mathrm{L}_{\mathrm{X,max}}=40.26\mathrel{\mathrm{erg}/\mathrm{s}}$).
Notably, escaper ULXs are less numerous, with a median count of $31$ compared to $176$ for in-cluster ULXs. They also typically have more evolved companion stars, with $65\%$ having CHeB donors compared to $55\%$ HG donors for in-cluster ULXs. The duty cycles and ULX phase onset times are similar for both populations.
The ejection processes preferentially remove less massive systems, explaining the lower masses observed in escaper ULXs. The higher fraction of neutron star accretors among escapers could be due to the higher natal kicks and retention of more massive black holes within the cluster potential.
These findings have important implications for understanding the origin and evolution of field ULXs. A significant fraction of observed field ULXs may have originated in globular clusters, which could explain some of the observed diversity in ULX populations. Furthermore, the differences between escaper and in-cluster ULXs provide valuable insights into how environment and dynamical history can influence the properties of these extreme systems.
In conclusion, our simulations reveal that escaper ULXs form a significant and distinct population with properties that differ from their in-cluster counterparts. These differences provide valuable insights into the formation and evolution of ULXs in various environments and highlight the need for comprehensive observational surveys to fully characterize the ULX population across different galactic settings.
4 Discussion
The results presented in this study are specific to the simulations employed and may not capture general trends applicable to all GC environments. The parameter choices were carefully tailored to conditions characteristic of GCs, informed by insights from analogous studies.
A key limitation in the broader study of ULXs in GCs is the paucity of confirmed observations within the Milky Way and its immediate vicinity. The substantial distances involved often complicate the detection and characterization of low-luminosity donor stars, which are believed to comprise a significant fraction of ULX donors (e.g. Wiktorowicz et al. 2017).
Furthermore, a critical challenge for this study lies in the difficulty of unambiguously differentiating GC-associated ULXs from those in the field environment. Observational uncertainties frequently obscure the distinction, as localization on the plane of a GC does not inherently confirm membership within the cluster. Determining the precise distance to the source is often fraught with uncertainties, leaving room for the possibility that the ULX is a foreground or background source, merely coincident with the GC in projection. This ambiguity highlights the importance of complementary approaches, such as proper motion studies or radial velocity measurements, to confirm GC membership conclusively.
4.1 Comparison with field populations
We find a high fraction of NS accretors among escapers ($40\%$) compared to in-cluster ULXs ($4\%$). This is similar to the predictions of Wiktorowicz et al. (2019), who found that NS ULXs outnumber BH ULXs in regions with constant star formation and solar metallicity for ages above $\sim 1\mathrel{\mathrm{Gyr}}$.
The properties of our escaper NS ULXs show some similarities with typical field NS ULXs described by Wiktorowicz et al. (2017). They found that field NS ULXs typically have $\sim 1.3\mathrel{\mathrm{M}_{\odot}}$ NS accretors and $\sim 1.0\mathrel{\mathrm{M}_{\odot}}$ Red Giant donors. This is comparable to our escaper ULXs with NS accretors, which have low-mass $\lesssim 2\mathrel{\mathrm{M}_{\odot}}$ evolved (HeMS) donors. Our systems are very compact ($a\lesssim 15\mathrel{\mathrm{R}_{\odot}}$) which may explain the lower fraction of giant donors.
Our escaper NS ULXs have a median maximum X-ray luminosity of $\log_{10}\mathrm{L}_{\mathrm{X,max}}=39.22\mathrel{\mathrm{erg}/\mathrm{s}}$, with a few BH ULXs exceeding $10^{40}\mathrel{\mathrm{erg}/\mathrm{s}}$. This stays in contrast to the field populations, where a significant fraction of more luminous sources is also expected (e.g., Wiktorowicz et al. 2015). In the case of globular clusters, we find that such extreme ULXs are more likely to be retained within the cluster.
[…]
The temporal distribution of ULXs in our simulations differs from that observed in field populations. Kuranov et al. (2021) found that the maximum number of ULXs ($\sim 10$ for a star formation rate of $10\mathrel{\mathrm{M}_{\odot}}/yr$) is reached $\sim 1\mathrel{\mathrm{Gyr}}$ after the beginning of star formation in field populations. Our escaper ULXs show a different temporal distribution with a more uniform appearance of escaper ULXs through the GC lifespan.
It is worth noting that even field populations are not entirely free from dynamical influences. Klencki et al. (2017) showed that wide binaries in the field are affected by multiple weak interactions (fly-bys) which can impact binary evolution. This suggests that the distinction between cluster-formed and field-formed ULXs may not be as clear-cut as previously thought, and that dynamical effects could play a role in shaping ULX populations across various environments.
4.2 Counterparts
The identification and characterization of ULX counterparts is crucial for understanding the nature of these systems, yet it remains a significant observational challenge (Heida et al. 2016; Wiktorowicz et al. 2021). While many ULXs are found in star-forming regions, their faintness in the optical and near-infrared (NIR), combined with the crowded environments, makes it difficult to confirm donor stars (López et al. 2020; Allak et al. 2022; Heida et al. 2019). The observed optical emission can be a combination of light from the accretion disc and/or the donor star (Heida et al. 2016). Accurate astrometry is essential for identifying counterparts, particularly for distinguishing between potential candidates within the error radii of ULX positions (Allak et al. 2022).
Red supergiants (RSGs) have also been identified as potential donors in ULXs (Heida et al. 2016, 2019; López et al. 2020). However, several sources initially classified as RSGs have been found to be nebulae, making the classification challenging (López et al. 2020). Actually, the fraction of ULXs with RSG donors is potentially higher than predicted by binary evolution models (López et al. 2020).
It’s important to note that observational biases exist that favour the detection of more luminous and massive companions, such as RSGs (e.g. Heida et al. 2019). Additionally, systems may be old, which makes it less likely for the donor to be a supergiant (Wiktorowicz et al. 2017). Therefore, ULX counterparts remain ’candidates’ and need more thorough survey searches (Wiktorowicz et al. 2021).
4.3 Expectations for Antennae galaxy
The correlation between X-ray source positions and stellar clusters has been previously noted in starburst galaxies (e.g. Kaaret et al. 2004). The study of the Antennae galaxies by Poutanen et al. (2013) reveals a highly significant association between bright X-ray sources and stellar clusters, although most of these X-ray sources are located outside of the clusters.
This finding is consistent with the idea that many ULXs are massive X-ray binaries that have been ejected from their birth clusters, rather than being intermediate mass black holes (IMBHs). Both ULXs and XRBs show a strong association with young stellar clusters, indicating a common origin.
The displacements of ULXs from centres of their neighbouring star clusters are statistically significant and range up to 300 parsecs. This is likely due to the ejection of massive binaries from the clusters.
The ejection mechanism suggested by Poutanen et al. (2013), supernova kicks and few-body encounters are consistent with the results of our work, where ULX among escapers have much higher NS accretor fraction and much smaller separations.
Poutanen et al. (2013) provides evidences that dynamical ejection can be a major formation mechanism for these systems in the field.
4.4 Future prospects
Future studies of ULXs in globular clusters should include more simulations to better estimate statistical effects, errors, and the significance of conclusions. A more consistent survey of the parameter space is needed, as well as the incorporation of evolutionary parameters (both stellar and binary evolution) beyond dynamical ones.
In the future work we plan to include the effect of beaming in ULXs (e.g. Lasota & King 2023). Beaming increases the apparent luminosity of ULXs, but decreases the probability of observation due to possible misalignment (Wiktorowicz et al. 2019; Khan et al. 2022). Population synthesis simulations, Wiktorowicz et al. (e.g. 2019) indicate that the majority of neutron star (NS) ULXs are beamed. The beaming factor is dependent on the mass transfer rate, with higher mass transfer rates resulting in stronger beaming (King 2009).
Additionally, the importance of wind Roche-lobe overflow (WRLOF; see e.g., Wiktorowicz et al. 2021; Zuo et al. 2021, and references therein) for NS ULXs with (super)giant donors should be explored. WRLOF can lead to stable mass transfer even with large mass ratios, and can boost mass transfer rates to reach ULX luminosity levels.
5 Conclusions
This study presents the first numerical investigation into the formation and evolution of ULXs within GCs, utilizing a subset of simulations to explore parameter dependencies.
Our simulations reveal that dynamical interactions may play a critical role in ULX formation, particularly in environments with high stellar density. Even if the initial binary fraction is low, the interactions can produce many ULXs. On average, we find that approximately $96\%$ of ULXs in our simulations have BH companions and the number can be even higher for very young clusters ($\lesssim 300\mathrel{\mathrm{Myr}}$). Among escaper, ULXs have a much higher fraction of NS accretors ($\sim 40\%$).
Our simulations show that the ratio of escaper ULXs to in-cluster ULXs is approximately 1:7, but nearly 2:1 for ULX with NS accretors. This ratio is the highest in tidally filling clusters, even with high initial binary fractions, where rapid ejections can hinder in-cluster ULX formation by ejecting the progenitors.
Our findings suggest that the relative scarcity of ULXs observed in GCs may be attributed to the advanced age of their stellar populations. In contrast, field populations usually have continuous star formation.
The apparent absence of ULXs in Milky Way GCs aligns more closely with our models of initially tidally filling clusters rather than non-tidally filling ones.
Furthermore, our results suggest that field populations may be significantly ”polluted” by ULXs ejected from GCs (escapers), which could contribute to the relative underrepresentation of ULXs in GCs compared to field environments. Escapers can have properties very similar to the field populations or form dinsinct configurations.
Future observational efforts should prioritize the comparison of ULX properties across diverse environments, including GCs and the galactic field. Such studies will be critical for validating our theoretical predictions and advancing our understanding of the diverse formation channels of ULXs.
Acknowledgments
GW, MG, AH, LH were supported by the Polish National Science Center (NCN) through the grant 2021/41/B/ST9/01191 AA acknowledges support for this paper from project No. 2021/43/P/ST9/03167 co-funded by the Polish National Science Center (NCN) and the European Union Framework Programme for Research and Innovation Horizon 2020 under the Marie Skłodowska-Curie grant agreement No. 945339. For the purpose of Open Access, the authors have applied for a CC-BY public copyright license to any Author Accepted Manuscript (AAM) version arising from this submission.
Data Availability
Data available on request.

### v3

@@
 Furthermore, the study identifies a population of ”escaper” ULXs, which originate in GCs but are ejected and emit X-rays outside the cluster.
+Our simulations reveal that these escapers constitute about one-seventh of the total ULX population.
+However, for neutron star accretors specifically, escapers are twice as common as their in-cluster counterparts.
+Notably, only $4\%$ of in-cluster ULXs contain neutron star accretors.
 These escapers may significantly contribute to the observed field ULX population.
@@
 Ultraluminous X-ray sources (ULXs) have been a subject of intense study in the field of high-energy astrophysics for several decades.
-These enigmatic objects, characterized by X-ray luminosities exceeding $10^{39}\mathrel{\mathrm{erg}/\mathrm{s}}$, challenge our understanding of accretion physics and compact object formation (see e.g.
-Kaaret et al.
-2017a, b, for a recent review).
-With over $1800$ known ULXs identified to date (Walton et al.
-2022), they represent a significant population of extreme X-ray emitters in the Universe, yet their fundamental nature remains a topic of ongoing debate.
-The diversity of ULXs has been highlighted in recent classification schemes (e.g.
-Wiktorowicz et al.
-2017), which attempt to categorize these objects based on their observational properties and potential underlying physical mechanisms.
-While the majority of ULXs were initially thought to harbor intermediate-mass black holes (IMBHs), our understanding has evolved considerably with the detection of pulsar accretors (Bachetti et al.
-2014).
+These enigmatic objects, characterized by X-ray luminosities exceeding $10^{39}\mathrel{\mathrm{erg}/\mathrm{s}}$, challenge our understanding of accretion physics and compact object formation (see e.g., Kaaret et al., 2017; King et al., 2023, for a recent review).
+With over $1800$ known ULXs identified to date (Walton et al., 2022), they represent a significant population of extreme X-ray emitters in the Universe, yet their fundamental nature remains a topic of ongoing debate.
+The diversity of ULXs has been highlighted in recent classification schemes (e.g., Wiktorowicz et al., 2017), which attempt to categorize these objects based on their observational properties and potential underlying physical mechanisms.
+While the majority of ULXs were initially thought to harbor intermediate-mass black holes (IMBHs), our understanding has evolved considerably with the detection of pulsar accretors (Bachetti et al., 2014).
 The importance of studying ULXs, particularly in the context of globular clusters (GC), cannot be overstated.
-These objects serve as laboratories for extreme accretion physics, potentially shedding light on the formation and evolution of double compact objects (DCOs) - the progenitors of gravitational wave sources.
-Moreover, ULXs continue to be considered as candidates for elusive IMBHs (Kaaret et al.
-2017a, and references therein).
-A critical aspect of ULX research involves identifying and characterizing their donor stars.
-Heida et al.
-(2016); López et al.
-(2020) reported that some ULXs had been identified with potential red super-gian (RSG) donors using infrared spectroscopy data.
-They examined a total of $113$ ULXs using photometric techniques in the infrared band, resulting in the identification of candidate counterparts for $38$ objects.
-Among these, the nature of $12$ sources was confirmed through spectroscopic analysis, with only five categorized as ULXs likely to have RSG donors.
-In this paper, we present a comprehensive study of ULXs in globular clusters, focusing on their stellar-mass manifestations.
+ULXs serve as laboratories for extreme accretion physics, potentially shedding light on the formation and evolution of double compact objects - the progenitors of gravitational wave sources.
+Moreover, ULXs continue to be considered as candidates for elusive IMBHs with masses in the range of $10^{2}$ to up to $10^{5}\mathrel{\mathrm{M}_{\odot}}$ (Kaaret et al., 2017, and references therein).
+The existence of IMBHs may provide crucial insights into the formation pathways of supermassive BHs observed at the centers of most galaxies.
+These intermediate-mass objects could also explain the presence of supermassive BHs observed at high redshift (Greene et al., 2020; Askar et al., 2023).
+In this paper, we present a comprehensive study of ULXs in GCs, focusing on their stellar-mass manifestations.
 The dense stellar environment introduces additional complexities and possibilities for exotic binary formations.
 By examining their properties, distribution, and potential formation mechanisms, we aim to contribute to the broader understanding of ULX physics and their role in compact object evolution.
-1.1 Globular Cluster environment
-GCs are dense, quasi-spherical collections of stars that orbit the centers of galaxies.
+Dense, quasi-spherical collections of stars known as GCs orbit the centers of galaxies.
 These typically old stellar systems, often containing millions of stars within a relatively small volume, are hubs of dynamical interactions.
-Their evolution is governed by the relaxation process connected with continuous distance gravitational interactions between all objects in the cluster.
-This leads to global phenomena like, core collapse, i.e.
-contraction of the cluster core and gradual increase in central density, and mass segregation, i.e.
-the tendency of heavier stars to occupy the central regions, while pushing the lighter stars outwards.
-Close dynamical interactions influence also the stars and binaries individually, influencing the binary parameters (e.g.
-the separation and eccentricity), or through exchanges (e.g.
-binary changes one of it’s component into another star), ejections (star or binary becomes unbound with the cluster), and binary formation (single stars became bound and form a binary).
-Such dynamical encounters in dense star clusters can also lead to the formation of close binary systems, including X-ray binaries (Pooley et al.
-2003; Ivanova et al.
-2005, 2008).
-Understanding how the fundamental dynamics influence the ULX formation is essential for interpreting the presence and characteristics of these exotic objects in GC environments.
-Figure 1: A schematic representation of the origins and eventual locations of ULX progenitors and ULXs in GCs.
+The masses of Milky Way GCs range from a few $10^{3}$ to a few $10^{6}\ \mathrel{\mathrm{M}_{\odot}}$, with half-mass radii typically between 1.5 pc and a few tens of pc, and ages spanning from 8 to 13.5 Gyr (Harris, 1996; Salaris & Weiss, 2002; VandenBerg et al., 2013; Baumgardt & Hilker, 2018).
+Detailed properties of specific Galactic GCs can be found in the catalogs of Harris (1996, updated 2010) [^1] https://physics.mcmaster.ca/~harris/mwgc.dat and Baumgardt & Hilker (2018) [^2] https://people.smp.uq.edu.au/HolgerBaumgardt/globular/parameter.html.
+The evolution of GCs is governed by the relaxation process connected with continuous distant gravitational interactions between all objects in the cluster.
+This leads to global phenomena like core collapse, that is contraction of the cluster core and gradual increase in central density, and mass segregation, that is the tendency of heavier stars to occupy the central regions while pushing lighter stars outwards.
+Close dynamical interactions also influence stars and binaries individually, affecting binary parameters (e.g., separation and eccentricity), or through exchanges (e.g., binary changes one of it’s component into another star), ejections (star or binary becomes unbound with the cluster), and binary formation (single stars became bound and form a binary).
+Such dynamical encounters in dense star clusters can lead to the formation of close binary systems, including X-ray binaries (Pooley et al., 2003; Ivanova et al., 2005, 2008; Kremer et al., 2018, 2019), cataclysmic variables (Oh et al., 2024; Belloni et al., 2019), and gravitational wave progenitors (e.g., Benacquista & Downing, 2013; Mandel & Broekgaarden, 2022, and references therein).
+Recent studies have emphasized the importance of GC dynamics in shaping exotic stellar populations (Di Carlo et al., 2021; Oh et al., 2024), highlighting the role of these environments in producing and evolving compact object binaries.
+While the majority of ULXs are detected in star-forming regions of galaxies, there is increasing evidence for their presence in GCs as well.
+These GCULXs are found exclusively in extragalactic GCs, all of which are unresolved at distances of ¿16 Mpc, with no ULXs detected in Milky Way GCs to date.
+Pioneering observations of GCULXs have been reported in several galaxies, with varied GC environments, spectral behavior, and variability patterns.
+Maccarone et al.
+(2007) discovered the first ULX in a GC (NGC 4472).
+It exhibits an X-ray luminosity of $4\times 10^{39}\mathrel{\mathrm{erg}/\mathrm{s}}$ and rapid variability.
+Similar detections were reported in NGC 1399 (Shih et al., 2010; Irwin et al., 2010), with more GCULXs in NGC 4472 (Maccarone et al., 2011) and NGC 4649 (Roberts et al., 2012).
+Some sources identified in GCs exhibited flares just above the ULX-defining limit of $10^{39}\mathrel{\mathrm{erg}/\mathrm{s}}$ (Irwin et al., 2016; Sivakoff et al., 2005).
+Recent observations have extended the known GCULX population, including seven additional GCULXs associated with M87 GCs (Dage et al., 2020), three in NGC 1316 (Dage et al., 2021), $10$ new candidates in GCs of massive ($>10^{11.5}M_{\odot}$) early-type galaxies (Thygesen et al., 2023), and two more in NGC 4261 (Nair et al., 2023).
+The current population of approximately 30 known GCULXs, identified through systematic searches of thousands of GC systems in elliptical galaxies, represents a remarkably rare phenomenon compared to the $\sim 1800$ ULXs identified in other galactic environments (Walton et al., 2022).
+In dense stellar systems, compact objects can be readily ejected, which can considerably impact the formation channels of such objects (e.g., Kulkarni et al., 1993).
+Furthermore, GCs typically exhibit high ages and, consequently, fewer higher-mass stars.
+Given their presence in older stellar populations, GCULXs are likely powered by low-mass X-ray binaries, in contrast to field ULXs which are typically associated with massive donor stars but often lack optical counterparts.
+Given their presence in older stellar populations, GCULXs are likely powered by low-mass X-ray binaries, in contrast to field ULXs which are expected to be also associated with low-mass as well as massive donor stars, but often lack optical counterparts.
+These characteristics contribute to the intriguing nature of GCULXs.
+2 Methodology
+2.1 MOCCA Monte Carlo code
+The MOCCA (MOnte Carlo Cluster evolution Code; Hypki & Giersz, 2013) is a state-of-the-art numerical tool that combines the Monte Carlo method for stellar dynamics (Hénon, 1971; Stodolkiewicz, 1986) with detailed stellar and binary evolution algorithms from the Binary Star Evolution code (BSE; Hurley et al., 2000, 2002) with recent updates (level-C from Kamlah et al., 2022, and references therein) and low-N scattering code FEWBODY (Fregeau et al., 2004) to follow close dynamical interactions.
+This hybrid approach makes MOCCA particularly well-suited for studying the formation and evolution of exotic objects like ULXs in dense stellar environments, offering an optimal balance between computational efficiency and physical accuracy.
+Recent improvements to the MOCCA code (Hypki et al., 2022, 2025; Giersz et al., 2024) have significantly enhanced its capabilities for ULX studies.
+First of all, the mass transfer calculations now incorporate a more sophisticated treatment of super-Eddington accretion, crucial for modeling ULX systems (Shakura & Sunyaev, 1973; Wiktorowicz et al., 2015).
+Additionally, improvements in core radii calculations improved the Roche-lobe overflow calculations.
+The introduction of detailed evolutionary tracking provide unprecedented insight into the spatial distribution and evolutionary pathways of ULX systems within clusters equivalent to detailed outputs from codes like startrack (Belczynski et al., 2008).
+A major advancement relevant to this study is MOCCA’s capability to handle multiple stellar populations, a phenomenon now recognized as ubiquitous in GCs (e.g., Piotto et al., 2015; Renzini et al., 2015; Gratton et al., 2019; Bastian & Lardo, 2018; Milone & Marino, 2022).
+While a comprehensive analysis of multiple stellar populations in our models is presented in Hypki et al.
+(2022, 2025); Giersz et al.
+(2024), here we specifically investigate their impact on ULX formation and evolution.
+The code now tracks different stellar populations with distinct chemical compositions, ages, and spatial distributions, allowing us to examine how properties specific to multiple stellar populations affect compact object formation and ULX characteristics.
+For detailed analysis of simulational results on multiple stellar populations and all recent changes to the MOCCA code see Giersz et al.
+(2024, and references therein).
+The X-ray luminosity ($\mathrm{L}_{\mathrm{X}}$) of accretion disks is calculated using the Shakura-Sunyaev model (Shakura & Sunyaev, 1973), accounting for both sub-Eddington and super-Eddington accretion regimes.
+The Eddington mass accretion rate is defined as $\dot{\mathrm{M}}_{\mathrm{Edd}}=2.2\times 10^{-8}\mathrm{M}_{\mathrm{acc}}\eta_{0.1}^{-1}\;[\mathrel{\mathrel{\mathrm{M}_{\odot}}\;\mathrm{yr}^{-1}}]$, where $\mathrm{M}_{\mathrm{acc}}$ is the accretor mass and $\eta\;(=10\eta_{0.1})$ is the accretion efficiency.
+We assumed $\eta_{0.1}=1$.
+The Eddington luminosity is then $\mathrm{L}_{\mathrm{Edd}}\approx 5.66\times 10^{45}\;\dot{\mathrm{M}}_{\mathrm{Edd}}\;[\text{erg s}^{-1}]$.
+The X-ray luminosity is calculated as:
+$\mathrm{L}_{\mathrm{X}}=\cases{\mathrm{L}_{\mathrm{Edd}}}(\dot{\mathrm{M}}/\dot{\mathrm{M}}_{\mathrm{Edd}}),&\text{for }\dot{\mathrm{M}}\leq\dot{\mathrm{M}}_{\mathrm{Edd}}\\ \mathrm{L}_{\mathrm{Edd}}[1+\ln(\dot{\mathrm{M}}/\dot{\mathrm{M}}_{\mathrm{Edd}})],\text{for }\dot{\mathrm{M}}>\dot{\mathrm{M}}_{\mathrm{Edd}}{}$ (1)
+where $\dot{\mathrm{M}}$ is the mass accretion rate. See Lasota & King (2023) for recent discussion of this formulation.
+2.2 Simulations
+Figure 1: Schematic representation of the origins and eventual locations of ULX progenitors and ULXs in GCs.
 Progenitors can either form and remain bound to the cluster, leading to in-cluster ULX emission (”IN-CLUSTER”), or be dynamically ejected and emit as ULXs outside the cluster (”ESCAPER”).
@@
 The key distinction lies in the dynamical imprint: ”ESCAPER” ULXs bear the hallmarks of GC interactions, unlike ”FIELD” systems, which provide a baseline for understanding the unique contribution of clusters to ULX formation.
-1.2 GCULX candidates
-While the majority of ULXs are detected in star-forming regions of galaxies, there is increasing evidence for their presence in GCs as well.
-Pioneering observations of GCULXs have been reported in several galaxies, with varied GC environments, spectral behavior, and variability patterns.
-Maccarone et al.
-(2007) discovered the first ULX in a GC (NGC 4472).
-It exhibits an X-ray luminosity of $4\times 10^{39}\mathrel{\mathrm{erg}/\mathrm{s}}$ and rapid variability.
-Similar detections were reported in NGC 1399 (Shih et al.
-2010; Irwin et al.
-2010), with more GCULXs in NGC 4472 (Maccarone et al.
-2011) and NGC 4649 (Roberts et al.
-2012).
-Some sources identified in GCs exhibited flares just above the ULX-defining limit of $10^{39}\mathrel{\mathrm{erg}/\mathrm{s}}$ (Irwin et al.
-2016; Sivakoff et al.
-2005).
-Recent observations have extended the known GCULX population, including seven additional GCULXs associated with M87 GCs (Dage et al.
-2020), three in NGC 1316 (Dage et al.
-2021), and $10$ new candidates in GCs of massive ($>10^{11.5}M_{\odot}$) early-type galaxies (Thygesen et al.
-2023).
-In dense stellar systems, compact objects can be readily ejected, which can considerably impact the formation channels of such objects (e.g.
-Kulkarni et al.
-1993).
-Furthermore, GCs typically exhibit high ages and, consequently, fewer higher-mass stars.
-These characteristics contribute to the intriguing nature of GCULXs.
-2 Methodology
-2.1 MOCCA Monte Carlo Code
-The MOCCA (MOnte Carlo Cluster evolution Code; Hypki & Giersz 2013) is a state-of-the-art numerical tool that combines the Monte Carlo method for stellar dynamics (Hénon 1971; Stodolkiewicz 1986) with detailed stellar and binary evolution algorithms from the BSE code (Hurley et al.
-2000, 2002) with recent updates (level-C from Kamlah et al.
-2022, and references therein) and low-N scattering code FEWBODY (Fregeau et al.
-2004) to follow close dynamical interactions.
-This hybrid approach makes MOCCA particularly well-suited for studying the formation and evolution of exotic objects like ULXs in dense stellar environments, offering an optimal balance between computational efficiency and physical accuracy.
-Recent improvements to the MOCCA code (Hypki et al.
-2022, 2025; Giersz et al.
-2024) have significantly enhanced its capabilities for ULX studies.
-First of all, the mass transfer calculations now incorporate a more sophisticated treatment of super-Eddington accretion, crucial for modeling ULX systems (Shakura & Sunyaev 1973; Wiktorowicz et al.
-2015).
-Additionally, improvements in core radii calculations improved the RLOF calculations.
-The introduction of detailed evolutionary tracking provide unprecedented insight into the spatial distribution and evolutionary pathways of ULX systems within clusters equivalent to detailed outputs from codes like startrack (Belczynski et al.
-2008).
-A major advancement relevant to this study is MOCCA’s capability to handle multiple stellar populations (MSPs), a phenomenon now recognized as ubiquitous in globular clusters (e.g.
-Piotto et al.
-2015; Renzini et al.
-2015; Gratton et al.
-2019; Bastian & Lardo 2018; Milone & Marino 2022).
-While a comprehensive analysis of MSPs in our models is presented in Hypki et al.
-(2022, 2025); Giersz et al.
-(2024), here we specifically investigate their impact on ULX formation and evolution.
-The code now tracks different stellar populations with distinct chemical compositions, ages, and spatial distributions, allowing us to examine how MSP-specific properties affect compact object formation and ULX characteristics.
-All recent changes to the MOCCA code are described in details in Giersz et al. (2024).
-2.2 Simulations
-The MOCCA code employs hundreds of parameters to govern the dynamical and stellar evolution of globular clusters.
-Based on previous studies of compact object populations - including cataclysmic variables (Belloni et al.
-2017), ULXs in the field (Wiktorowicz et al.
-2019), and double white dwarf systems (Hellström et al.
-2024) - we identified key parameters that most significantly influence ULX formation and evolution.
+The MOCCA code employs hundreds of parameters to govern the dynamical and stellar evolution of GCs.
+Based on previous studies of compact object populations - including cataclysmic variables (Belloni et al., 2017), ULXs in the field (Wiktorowicz et al., 2019), and double white dwarf systems (Hellström et al., 2024) - we identified key parameters that most significantly influence ULX formation and evolution.
 Our parameter selection strategy focused on those that control dynamics and initial stellar population properties.
@@
 For all simulations with dynamics enabled, we tracked the evolution of escaped systems to analyze the properties of their populations and potential ULXs that can form after their progenitors are ejected from the cluster.
-These escapers evolve as isolated binaries, similar to field binary evolution (e.g., Fragos et al.
-2015; Wiktorowicz et al.
-2019; Zuo et al.
-2021).
-Systems can escape the cluster through multiple mechanisms: (1) dynamical interactions, particularly strong few-body encounters, (2) natal kicks from supernovae explosions, including both direct and Blaauw kicks, (3) gradual two-body relaxation (Spitzer & Hart 1971), and (4) tidal stripping of outer cluster regions by the host galaxy’s gravitational field (Gnedin & Ostriker 1997).
+These escapers evolve as isolated binaries, similar to field binary evolution (e.g., Fragos et al., 2015; Wiktorowicz et al., 2019; Zuo et al., 2021).
+Systems can escape the cluster through multiple mechanisms: (1) dynamical interactions, particularly strong few-body encounters, (2) natal kicks from supernovae explosions, including both direct and Blaauw kicks, (3) gradual two-body relaxation (Spitzer & Hart, 1971), and (4) tidal stripping of outer cluster regions by the host galaxy’s gravitational field (Gnedin & Ostriker, 1997).
 The galactocentric distance ($\mathrm{R}_{\mathrm{gc}}$) for each simulation was scaled to maintain an initial tidal radius of $\mathrm{R}_{\mathrm{tid}}\approx 43\mathrel{\mathrm{pc}}$ across all models.
 This normalization of the tidal radius facilitates meaningful comparisons between simulations, particularly when analyzing the degree of tidal filling and cluster structural parameters.
-The tidal radius follows the relation $\mathrm{R}_{\mathrm{tid}}\propto\mathrm{R}_{\mathrm{gc}}^{2/3}\mathrm{M}_{\mathrm{tot}}^{1/3}$ (e.g.
-Webb et al.
-2013), where $\mathrm{M}_{\mathrm{tot}}$ is the cluster mass.
+The tidal radius follows the relation $\mathrm{R}_{\mathrm{tid}}\propto\mathrm{R}_{\mathrm{gc}}^{2/3}\mathrm{M}_{\mathrm{tot}}^{1/3}$ (e.g., Webb et al., 2013), where $\mathrm{M}_{\mathrm{tot}}$ is the cluster mass.
 This scaling ensures that clusters experience comparable relative tidal forces despite different masses and orbital parameters.
-We note that while the initial tidal radii are identical, the clusters’ subsequent evolution may lead to different filling factors111filling factor is R_tid/R_h, which is initially $\sim 43$ for non-tidally filling simulations and $\sim 3.6$ for tidally filling ones.
+We note that while the initial tidal radii are identical, the clusters’ subsequent evolution may lead to different filling factors333filling factor is $\mathrm{R}_{\mathrm{tid}}$/$\mathrm{R}_{\mathrm{h}}$, which is initially $\sim 43$ for non-tidally filling simulations and $\sim 3.6$ for tidally filling ones.
 depending on their internal dynamics and mass-loss history.
-Table 1: Model parameters and initial properties
-Label $\mathrm{N}_{\mathrm{pop}}$ $\mathrm{R}_{\mathrm{h,2}}/\mathrm{R}_{\mathrm{h,1}}$ $\mathrm{f}_{\mathrm{bin}}$ $\mathrm{R}_{\mathrm{gc}}$ $\mathrm{R}_{\mathrm{h}}$ $\mathrm{M}_{\mathrm{tot}}/10^{6}$ $\mathrm{R}_{\mathrm{c}}$ $\mathrm{M}_{\mathrm{c}}/10^{3}$ $\log_{10}(\rho_{\mathrm{c}})$ $\mathrm{t}_{\mathrm{h,rel}}$
-npop1-fb10-TF 1 10% $2.01$ $11.90$ $0.69$ $5.85$ $89.10$ $2.64$ $10,888.00$
-npop1-fb10-nTF 1 10% $2.01$ $1.00$ $0.69$ $0.49$ $89.10$ $5.87$ $265.22$
-npop1-fb95-TF 1 95% $1.96$ $11.93$ $0.72$ $5.81$ $93.93$ $2.61$ $6,386.70$
-npop1-fb95-nTF 1 95% $1.96$ $1.00$ $0.72$ $0.49$ $93.93$ $5.84$ $155.06$
-npop2-cpop05-fb10-TF 2 0.05 10% $2.01$ $11.69$ $0.66$ $0.13$ $12.71$ $356$,505.306505.306,505.306 , 505.30
-npop2-cpop05-fb10-TF-NF 2 0.05 10% $2.40$ $11.81$ $0.46$ $6.26$ $68.94$ $2.41$ $9,157.70$
-npop2-cpop05-fb10-nTF 2 0.05 10% $2.01$ $0.99$ $0.66$ $0.01$ $12.71$ $9.57$ $160.33$
-npop2-cpop05-fb10-nTF-NF 2 0.05 10% $2.40$ $1.00$ $0.46$ $0.53$ $68.94$ $5.63$ $225.73$
-npop2-cpop05-fb95-TF 2 0.05 95% $1.96$ $11.75$ $0.69$ $0.12$ $13.03$ $6.37$ $3,790.60$
-npop2-cpop05-fb95-TF-NF 2 0.05 95% $2.35$ $11.77$ $0.48$ $6.18$ $71.69$ $2.39$ $5,342.70$
-npop2-cpop05-fb95-nTF 2 0.05 95% $1.96$ $1.00$ $0.69$ $0.01$ $13.03$ $9.59$ $93.85$
-npop2-cpop05-fb95-nTF-NF 2 0.05 95% $2.35$ $1.00$ $0.48$ $0.53$ $71.69$ $5.61$ $132.31$
-npop2-cpop2-fb10-TF 2 0.2 10% $2.01$ $11.80$ $0.66$ $0.50$ $12.84$ $4.55$ $7,083.50$
-npop2-cpop2-fb10-TF-NF 2 0.2 10% $2.40$ $11.81$ $0.46$ $6.09$ $64.82$ $2.40$ $9,157.10$
-npop2-cpop2-fb10-nTF 2 0.2 10% $2.01$ $1.00$ $0.66$ $0.04$ $12.84$ $7.77$ $174.58$
-npop2-cpop2-fb10-nTF-NF 2 0.2 10% $2.40$ $1.00$ $0.46$ $0.52$ $64.82$ $5.61$ $225.72$
-npop2-cpop2-fb95-TF 2 0.2 95% $1.96$ $11.77$ $0.69$ $0.49$ $13.12$ $564$,079.504079.504,079.504 , 079.50
-npop2-cpop2-fb95-TF-NF 2 0.2 95% $2.35$ $11.77$ $0.48$ $6.12$ $70.32$ $2.35$ $5,342.50$
-npop2-cpop2-fb95-nTF 2 0.2 95% $1.96$ $1.00$ $0.69$ $0.04$ $13.12$ $7.77$ $101.01$
-npop2-cpop2-fb95-nTF-NF 2 0.2 95% $2.35$ $1.00$ $0.48$ $0.52$ $70.32$ $5.56$ $132.31$
-222 List of simulations with differentiating parameters and some initial properties (at $t=0$).
-Label identifies particular simulation and will be used throughout the paper.
-TF/nTF in the label means initially tidally filling or non-tidally filling.
-NF, when present, means that new features were included in this model (see text for detailes); $\mathrm{N}_{\mathrm{pop}}$ number of stellar generations; $\mathrm{R}_{\mathrm{h,2}}/\mathrm{R}_{\mathrm{h,1}}$ ratio between the $\mathrm{R}_{\mathrm{h}}$ of the second and first stellar generation; $\mathrm{f}_{\mathrm{bin}}$ binary fraction; $\mathrm{R}_{\mathrm{gc}}$ galactocentric radius [$\mathrel{\mathrm{kpc}}$]; $\mathrm{R}_{\mathrm{h}}$ half-mass radius for the first population [$\mathrel{\mathrm{pc}}$]; $\mathrm{M}_{\mathrm{tot}}$ total stellar mass [$\mathrel{\mathrm{M}_{\odot}}$]; $\mathrm{R}_{\mathrm{c}}$ core radius [$\mathrel{\mathrm{pc}}$]; $\mathrm{M}_{\mathrm{c}}$ core mass [$\mathrel{\mathrm{M}_{\odot}}$]; $\rho_{\mathrm{c}}$ central density [$\mathrel{\mathrm{M}_{\odot}}/\mathrel{\mathrm{pc}}^{3}$]; $\mathrm{t}_{\mathrm{h,rel}}$ Spitzer half-mass relaxation time [$\mathrel{\mathrm{M}_{\odot}}$].
-All simulations include the escapers information (see text).
-For all these simulations the counterpart field population (with dynamics turned off) was calculated (nodyn in the label; not shown in the table) with the same parameters.
 We performed a comprehensive set of simulations exploring different initial conditions and physical parameters.
@@
 Some simulations incorporate the new features (NF) described below, particularly relevant for multiple population scenarios.
-For each simulation, we provide key structural parameters at $t=0\mathrel{\mathrm{Myr}}$, including the galactocentric radius ($\mathrm{R}_{\mathrm{gc}}$), half-mass radius ($\mathrm{R}_{\mathrm{h}}$), total stellar mass ($\mathrm{M}_{\mathrm{tot}}$), core radius ($\mathrm{R}_{\mathrm{c}}$; according to Casertano & Hut 1985), core mass ($\mathrm{M}_{\mathrm{c}}$, i.e.
-mass inside $\mathrm{R}_{\mathrm{c}}$), and central density ($\rho_{\mathrm{c}}$; according to Casertano & Hut 1985).
+For each simulation, we provide key structural parameters at $t=0\mathrel{\mathrm{Myr}}$, including the galactocentric radius ($\mathrm{R}_{\mathrm{gc}}$), half-mass radius ($\mathrm{R}_{\mathrm{h}}$), total stellar mass ($\mathrm{M}_{\mathrm{tot}}$), core radius ($\mathrm{R}_{\mathrm{c}}$; according to Casertano & Hut, 1985), core mass ($\mathrm{M}_{\mathrm{c}}$, i.e., mass inside $\mathrm{R}_{\mathrm{c}}$), and central density ($\rho_{\mathrm{c}}$; according to Casertano & Hut, 1985).
 Additionally, we computed corresponding field populations with dynamics turned off (denoted by nodyn in the model labels) to serve as control cases, though these are not shown separately in the table.
 To better compare simulations with different binary fractions, we made several adjustments to the parameter space.
-Since a lower binary fraction results in lower total stellar mass (as binary systems typically have higher masses than single stars), we adjusted the initial number of objects to have similar half-mass relaxation time for both models: simulations with $\mathrm{f}_{\mathrm{bin}}=95\%$ used $n=600,000$ objects, while those with $\mathrm{f}_{\mathrm{bin}}=10\%$ used $n=1,063,635$ objects.We also modified the initial distribution of semi-major axes.
-For simulations with $\mathrm{f}_{\mathrm{bin}}=95\%$, we used the modified version of the Kroupa (1995) period distribution (see Belloni et al.
-2017) for stars with $M<5\mathrel{\mathrm{M}_{\odot}}$, and the distributions from Sana et al.
+Since a lower binary fraction results in lower total stellar mass (as binary systems typically have higher masses than single stars), we adjusted the initial number of objects to have a similar half-mass relaxation time for both models: simulations with $\mathrm{f}_{\mathrm{bin}}=95\%$ used $n=600,000$ objects, while those with $\mathrm{f}_{\mathrm{bin}}=10\%$ used $n=1,063,635$ objects.
+We also modified the initial distribution of semi-major axes.
+For simulations with $\mathrm{f}_{\mathrm{bin}}=95\%$, we used the modified version of the Kroupa (1995) period distribution (see Belloni et al., 2017) for stars with $M<5\mathrel{\mathrm{M}_{\odot}}$, and the distributions from Sana et al.
 (2012); Oh et al.
@@
 (2012) period distribution for $M>5\mathrel{\mathrm{M}_{\odot}}$.
-Additionally, for the low binary fraction simulations, we disabled eigenevolution (Belloni et al.
-2017).
+Additionally, for the low binary fraction simulations, we disabled eigenevolution (Belloni et al., 2017).
 The separation was limited to the maximal value of $100\mathrel{\mathrm{AU}}$
-In a part of our simulations (marked with NF in label), we introduced several modifications to investigate the influence of multiple stellar populations (see Giersz et al.
-2024, for details).
+In a part of our simulations (marked with NF in label), we introduced NF to investigate the influence of multiple stellar populations (see Giersz et al., 2024, for details).
 Specifically, we implemented a delay time ($t_{\rm delay}$) of $100\mathrel{\mathrm{Myr}}$ for the second population, during which these stars act as gas particles without undergoing stellar evolution, relaxation, or dynamical interactions.
@@
 To accommodate this, we activated AGB wind accumulation for the second population formation.
-The cluster’s orbital evolution was modified by implementing an artificial orbit change at $t=1000\mathrel{\mathrm{Myr}}$, where the galactocentric distance increases by a factor of 2 (see Giersz et al.
-2024, for details).
+The cluster’s orbital evolution was modified by implementing an artificial orbit change at $t=1000\mathrel{\mathrm{Myr}}$, where the galactocentric distance increases by a factor of 2 (see Giersz et al., 2024, for details).
 Finally, we employed single-population scaling, where the scaling parameters are applied exclusively to the first population, while the second population’s velocities are independently normalized.
@@
 Except dynamical evolution parameters, our simulations incorporate several key prescriptions for stars and binaries, some of them particularly relevant to ULX formation, specifically: We employed the rapid supernova mechanism as described by Fryer et al.
-(2012) for both neutron star and black hole formation; Neutron star and black hole natal kicks were drawn from a Maxwellian distribution with $\sigma=265\mathrel{\mathrm{km}/\mathrm{s}}$ (Hobbs et al.
-2005); For black holes, these kicks were scaled by the fallback factor, which depends on the supernova prescription; Neutron stars formed through electron-capture supernovae were assigned substantially lower kicks with $\sigma=3\mathrel{\mathrm{km}/\mathrm{s}}$; Common envelope evolution was modeled using the BSE parameters $\alpha=0.5$ and $\lambda=0.0$.
-All simulated models in this study were set to a fixed metallicity of $\mathrm{Z}=0.001$.
-More details on the MOCCA parameters can be found in Kamlah et al. (2022); Hypki et al. (2022, 2025); Giersz et al. (2024).
+(2012) for both neutron star (NS) and black hole (BH) formation; Neutron star and BH natal kicks were drawn from a Maxwellian distribution with $\sigma=265\mathrel{\mathrm{km}/\mathrm{s}}$ (Hobbs et al., 2005); For BHs, these kicks were scaled by the fallback factor, which depends on the supernova prescription; Neutron stars formed through electron-capture supernovae were assigned substantially lower kicks with $\sigma=3\mathrel{\mathrm{km}/\mathrm{s}}$; Common envelope evolution was modeled using the parameters from the BSE code: $\alpha=0.5$ and $\lambda=0.0$.
+All simulated models in this study were set to a fixed metallicity of $\mathrm{Z}=0.001$, which represents the typical metallicity of Galactic globular clusters444see the Harris catalog of Milky Way globular clusters https://physics.mcmaster.ca/~harris/mwgc.dat (Harris, 1996, 2010 update)..
 3 Results
+ULXs in this study are defined as binary systems with neutron star or black hole accretors exhibiting X-ray luminosities exceeding $10^{39}\text{ erg s}^{-1}$.
+White dwarf-powered ULXs are excluded from this analysis.
 3.1 Number evolution of ULXs
-Figure 2: Evolution of ULX numbers in globular clusters.
+Figure 2: Evolution of ULX numbers in GCs.
 Expected number of ULXs ($\mathbb{E}[\mathcal{N}_{\mathrm{ULX}}]$) as a function of time since GC formation.
 All simulations with non-zero expected rates are presented.
-The line indicates the number of simulations with non-zero predictions (i.e.
-number of dots in this time bin).
+The line indicates the number of simulations with non-zero predictions (i.e., number of dots in this time bin).
 Each time bin spans $300\;\mathrel{\mathrm{Myr}}$ and the points are centered on these bins.
@@
 This trend persists across all simulations, irrespective of parameter variations.
-These findings align with previous studies on field populations, where dynamical interactions are negligible, which reported that the ULX phase predominantly occurs in the early evolutionary stages of binaries (e.g., Wiktorowicz et al.
-2017).
-Observational evidence also supports this conclusion (e.g.
-Wolter et al.
-2018).
+These findings align with previous studies on field populations, where dynamical interactions are negligible, which reported that the ULX phase predominantly occurs in the early evolutionary stages of binaries (e.g., Wiktorowicz et al., 2017).
+Observational evidence also supports this conclusion (e.g., Wolter et al., 2018).
 This result suggests either:
@@
   • Insufficient time has elapsed for dynamics to significantly influence ULX formation.
-The number of simulations with $\mathbb{E}[\mathcal{N}_{\mathrm{ULX}}]\neq 0$ decreases over time, resembling predictions for populations formed in burst-like star formation episodes (e.g.
-Wiktorowicz et al.
-2017).
+The number of simulations with $\mathbb{E}[\mathcal{N}_{\mathrm{ULX}}]\neq 0$ decreases over time, resembling predictions for populations formed in burst-like star formation episodes (e.g., Wiktorowicz et al., 2017).
 In GCs, this trend may result from a combination of system age effects on ULX formation (as observed in field populations) and the interplay of positive and negative effects related to dynamics.
-Generally, our results indicate a higher likelihood of observing ULXs in younger stellar clusters compared to older ones.
+Generally, our results indicate a higher likelihood of observing ULXs in younger stellar clusters ($\lesssim 300\mathrel{\mathrm{Myr}}$) compared to older ones, which is consistent with observational results (e.g., Dage et al., 2025).
 […]
@@
 3.2 Temporal evolution and parameter dependence of $\mathbb{E}[\mathcal{N}_{\mathrm{ULX}}]$
-Figure 3: Comprehensive overview of the $\mathbb{E}[\mathcal{N}_{\mathrm{ULX}}]$ evolution since the formation of a globular cluster (GC).
+Figure 3: Comprehensive overview of the $\mathbb{E}[\mathcal{N}_{\mathrm{ULX}}]$ evolution since the formation of a GC.
 Each panel represents a set of simulations with identical parameters: number of populations ($\mathrm{N}_{\mathrm{pop}}$), concentration ($\mathrm{R}_{\mathrm{h,2}}/\mathrm{R}_{\mathrm{h,1}}$), binary fraction ($\mathrm{f}_{\mathrm{bin}}$), and the presence of new features (NF).
@@
 Conversely, the binary fraction influences field populations, with a positive correlation between initial binaries and ULX progenitors, as there are no external processes to destroy ULX progenitors or create new ones in isolation.
-The incorporation of new features (NF) has a noticeable effect on two-population simulations with moderate concentration ($\mathrm{R}_{\mathrm{h,2}}/\mathrm{R}_{\mathrm{h,1}}=0.2$; Figure 3, middle column).
-Simulations with NF exhibit higher $\mathbb{E}[\mathcal{N}_{\mathrm{ULX}}]$ values compared to those without NF (nNF).
+The incorporation of NF has a noticeable effect on two-population simulations with moderate concentration ($\mathrm{R}_{\mathrm{h,2}}/\mathrm{R}_{\mathrm{h,1}}=0.2$; Figure 3, middle column).
+Simulations with NF exhibit higher $\mathbb{E}[\mathcal{N}_{\mathrm{ULX}}]$ values compared to those without NF.
 The NF do not affect field populations as they primarily relate to the time delay for the second population.
-The escaper populations generally align with the ULX population formed in situ, suggesting ongoing ejection (relacation and interactions), and thus a continuous supply of ULX progenitors to the escaper population, which later become ULXs while unbound to the cluster.
+The escaper populations generally align with the ULX population formed in situ, suggesting ongoing ejection (relaxation and interactions), and thus a continuous supply of ULX progenitors to the escaper population, which later become ULXs while unbound to the cluster.
 There are generally fewer ULXs among escapers than in in-cluster populations.
@@
 An exception occurs in simulations with relatively strong ULX formation ($\mathrm{N}_{\mathrm{pop}}=2$, $\mathrm{R}_{\mathrm{h,2}}/\mathrm{R}_{\mathrm{h,1}}=0.2$), where ULXs from escapers form at evolutionary ages not covered by field populations.
-For tidally filling clusters, the rate of ULX formation from escapers can exceed that of in-cluster formation, especially in simulations without new features (nNF).
+For tidally filling clusters, the rate of ULX formation from escapers can exceed that of in-cluster formation, especially in simulations without NF.
 3.3 Formation of ULXs
 This section focuses on ULX progenitors - systems that will evolve into ULXs - and their evolution up to the initial ULX phase.
-It is important to note that some systems may experience multiple ULX phases interspersed with periods of quiescence or XRB phases ($\mathrm{L}_{\mathrm{X}}<10^{39}\mathrel{\mathrm{erg}/\mathrm{s}}$).
+It is important to note that some systems may experience multiple ULX phases interspersed with periods of quiescence or X-ray binary phases ($\mathrm{L}_{\mathrm{X}}<10^{39}\mathrel{\mathrm{erg}/\mathrm{s}}$).
+These phase transitions typically occur on timescales of at least hundreds of years, driven by dynamical interactions or stellar evolution processes.
 Additionally, individual stars can participate in multiple ULXs, each with a different companion.
 This phenomenon is particularly prevalent for IMBHs (see Sect.
-3.4.1).
+3.5).
+While our models focus on these long-term changes, it’s worth noting that observations have shown both globular cluster and field ULXs can exhibit significant short-term variability on timescales of days to months (e.g., Maccarone et al., 2007; Shih et al., 2010; Brightman et al., 2023).
 Table 1: Progenitor properties: ZAMS progenitors, ULX onset time, and total ULXs for different configurations.
@@
 n-D $91.36\pm 174.29\;(12)$ $3.01\pm 2.27\;(13)$
-444 Number of progenitors on zams ($\mathrm{N}_{\mathrm{prog}}$), median fo the time when the ULX phase commences ($\mathrm{med}(\mathrm{t}_{\mathrm{ULX,start}})$), and the total (i.e.
-through the entire history) expected number of ULXs ($\mathbb{E}[\mathcal{N}_{\mathrm{ULX}}]_{total}$).
+666 Number of progenitors on ZAMS ($\mathrm{N}_{\mathrm{prog}}$), median of the time when the ULX phase commences ($\mathrm{med}(\mathrm{t}_{\mathrm{ULX,start}})$), and the total (i.e., through the entire history) expected number of ULXs ($\mathbb{E}[\mathcal{N}_{\mathrm{ULX}}]_{total}$).
 Number in parenthesis inform about the number of simulations with non-zero $\mathbb{E}[\mathcal{N}_{\mathrm{ULX}}]$ of specific configuration.
@@
 p stands for pristine and means that there were no exchanges in a binary before ULX formation.
-First symbol represents the primary (i.e.
-more massive star on ZAMS), whereas the second letter the secondary.
+First symbol represents the primary (i.e., more massive star on zero-age main sequence, or ZAMS), whereas the second letter the secondary.
 Results are divided into First and Second Populations and also between binaries and single stars.
 Values represent the mean and one sigma error for simulations in which the specific ULX where present.
-Table 2 provides a comprehensive overview of ULX progenitors on ZAMS, i.e.
-stars that later became components of ULXs.
-The most common progenitors are those of pristine ULXs, i.e.
-binaries that do not undergo exchanges or disruptions prior to the ULX phase.
+Table 1 provides a comprehensive overview of ULX progenitors on zero-age main sequence (ZAMS), that is stars that later became components of ULXs.
+The most common progenitors are those of pristine ULXs, that is binaries that do not undergo exchanges or disruptions prior to the ULX phase, although still they can dynamically interact and change their properties.
 In these systems, the primary typically becomes the accretor and the secondary the donor (A-D(p)).
@@
 However, stars born in binaries more frequently become components of ULXs, even in non-pristine cases.
-These systems invariably undergo significant interactions, such as exchanges or disruptions, before forming a ULX.
+These systems invariably undergo significant interactions, such as binary formation, exchanges or disruptions, before forming a ULX.
 While some of these ULXs can form relatively early ($\sim 100\mathrel{\mathrm{Myr}}$), the majority emerge in later evolutionary phases ($>3\mathrel{\mathrm{Gyr}}$).
@@
 3.4 ULX properties
-In this section we discuss the properties of in-cluster ULXs, i.e. these which reside inside the cluster. For discussion of ULXs among escapers see Section 3.7.
+In this section we discuss the properties of in-cluster ULXs, that is these which reside inside the cluster. For a discussion of ULXs among escapers see Section 3.8.
 Table 2: Properties of ULXs by group
@@
 777 Median properties and $95\%$ confidence intervals for ULX groups identified in this study.
-Column definitions: $\mathrm{k}_{\mathrm{acc}}$ accretor type; $\mathrm{k}_{\mathrm{don}}$ donor type; $\mathrm{M}_{\mathrm{acc}}$ accretor mass ($\mathrel{\mathrm{M}_{\odot}}$); $\mathrm{M}_{\mathrm{don}}$ donor mass ($\mathrel{\mathrm{M}_{\odot}}$); $a$ orbital separation ($\mathrel{\mathrm{R}_{\odot}}$); $e$ eccentricity; $\mathrm{L}_{\mathrm{X,max}}$ maximum X-ray luminosity ($\mathrel{\mathrm{erg}/\mathrm{s}}$); $\Delta\mathrm{t}_{\mathrm{ULX}}$ ULX lifetime ($\mathrel{\mathrm{Myr}}$); $\mathrm{t}_{\mathrm{phys,min}}$ ULX phase onset age.
+Column definitions: $\mathrm{k}_{\mathrm{acc}}$ accretor type (BH - black hole, NS - neutron star); $\mathrm{k}_{\mathrm{don}}$ donor type; $\mathrm{M}_{\mathrm{acc}}$ accretor mass ($\mathrel{\mathrm{M}_{\odot}}$); $\mathrm{M}_{\mathrm{don}}$ donor mass ($\mathrel{\mathrm{M}_{\odot}}$); $a$ orbital separation ($\mathrel{\mathrm{R}_{\odot}}$); $e$ eccentricity; $\mathrm{L}_{\mathrm{X,max}}$ maximum X-ray luminosity ($\mathrel{\mathrm{erg}/\mathrm{s}}$); $\Delta\mathrm{t}_{\mathrm{ULX}}$ ULX lifetime ($\mathrel{\mathrm{Myr}}$); $\mathrm{t}_{\mathrm{phys,min}}$ ULX phase onset age ($\mathrel{\mathrm{Myr}}$).
 Donor types: MS - main sequence, HG - Hertzsprung gap, RG - red giant, CHeB - core helium burning.
@@
 […]
-NS accretors in ULXs ($G_{\textrm{NS}}$) have median masses around $1.27\mathrel{\mathrm{M}_{\odot}}$.
+Within ULXs, NS accretors ($G_{\textrm{NS}}$) have median masses around $1.27\mathrel{\mathrm{M}_{\odot}}$.
 These systems typically have lower X-ray luminosities ($\log_{10}\mathrm{L}_{\mathrm{X,max}}\approx 39.22$) compared to their BH counterparts, but still clearly exceed the Eddington limit for a typical NS.
@@
 This variability in duty cycles may explain the transient nature observed in many ULXs and has implications for their detectability (cf.
-Wiktorowicz et al.
-2017).
+Wiktorowicz et al., 2017).
 A notable subset of our sample consists of IMBH-powered systems ($G_{\textrm{IMBH}}$).
 These ULXs are characterized by extremely high accretor masses (median $9351\mathrel{\mathrm{M}_{\odot}}$) and the highest X-ray luminosities in our sample ($\log_{10}\mathrm{L}_{\mathrm{X,max}}\approx 40.84$).
-Such luminous systems, called extreme-ULXs (Wiktorowicz et al.
-2019), are one of the main observational candidates for IMBHs and can provide insights into their formation and properties.
+Such luminous systems, called extreme-ULXs (Wiktorowicz et al., 2019), are one of the main observational candidates for IMBHs and can provide insights into their formation and properties.
 The diversity in donor types across ULX groups highlights the various evolutionary pathways that can lead to ULX formation.
-MS, HG, and RG donors are all well-represented, suggesting that ULXs can form and persist across a wide range of stellar evolutionary stages
-Table 4: Fraction of ULXs in Different Groups Across Simulations
-label $G_{\textrm{BH,MS}}$ $G_{\textrm{BH,HG}}$ $G_{\textrm{BH,CHeB}}$ $G_{\textrm{IMBH}}$ $G_{\textrm{NS}}$ other
-npop1-fb10-TF $0.17$ $0.33$ $0.36$ $0.11$ $0.04$
-npop1-fb10-nTF $0.16$ $0.32$ $0.30$ $0.12$ $0.11$
-npop1-fb10-nTF-nodyn $0.18$ $0.36$ $0.37$ $0.05$ $0.04$
-npop1-fb95-TF $0.11$ $0.40$ $0.33$ $0.12$ $0.04$
-npop1-fb95-nTF $0.12$ $0.39$ $0.28$ $0.11$ $0.11$
-npop1-fb95-nTF-nodyn $0.11$ $0.43$ $0.31$ $0.11$ $0.04$
-npop2-cpop05-fb10-TF $0.09$ $0.39$ $0.29$ $0.09$ $0.13$
-npop2-cpop05-fb10-nTF $0.07$ $0.29$ $0.20$ $0.33$ $0.04$ $0.06$
-npop2-cpop05-fb10-nTF-nodyn $0.10$ $0.46$ $0.32$ $0.04$ $0.07$
-npop2-cpop05-fb10-TF-NF $0.15$ $0.38$ $0.27$ $0.05$ $0.14$
-npop2-cpop05-fb10-nTF-NF $0.06$ $0.25$ $0.16$ $0.44$ $0.05$ $0.03$
-npop2-cpop05-fb10-nTF-nodyn-NF $0.11$ $0.48$ $0.33$ $0.08$
-npop2-cpop05-fb95-TF $0.11$ $0.29$ $0.42$ $0.11$ $0.06$
-npop2-cpop05-fb95-nTF $0.11$ $0.25$ $0.35$ $0.16$ $0.07$ $0.06$
-npop2-cpop05-fb95-nTF-nodyn $0.12$ $0.29$ $0.41$ $0.14$ $0.04$
-npop2-cpop05-fb95-TF-NF $0.13$ $0.27$ $0.44$ $0.10$ $0.06$
-npop2-cpop05-fb95-nTF-NF $0.10$ $0.18$ $0.29$ $0.34$ $0.05$ $0.04$
-npop2-cpop05-fb95-nTF-nodyn-NF $0.12$ $0.30$ $0.43$ $0.12$ $0.04$
-npop2-cpop2-fb10-TF $0.10$ $0.41$ $0.30$ $0.09$ $0.10$
-npop2-cpop2-fb10-nTF $0.06$ $0.28$ $0.19$ $0.36$ $0.04$ $0.07$
-npop2-cpop2-fb10-nTF-nodyn $0.10$ $0.46$ $0.32$ $0.04$ $0.07$
-npop2-cpop2-fb10-TF-NF $0.12$ $0.39$ $0.30$ $0.09$ $0.09$
-npop2-cpop2-fb10-nTF-NF $0.16$ $0.16$ $0.09$ $0.41$ $0.18$
-npop2-cpop2-fb10-nTF-nodyn-NF $0.11$ $0.48$ $0.33$ $0.08$
-npop2-cpop2-fb95-TF $0.13$ $0.26$ $0.44$ $0.13$ $0.05$
-npop2-cpop2-fb95-nTF $0.10$ $0.24$ $0.31$ $0.20$ $0.09$ $0.06$
-npop2-cpop2-fb95-nTF-nodyn $0.12$ $0.29$ $0.42$ $0.13$ $0.04$
-npop2-cpop2-fb95-TF-NF $0.12$ $0.28$ $0.45$ $0.11$ $0.05$
-npop2-cpop2-fb95-nTF-NF $0.13$ $0.17$ $0.22$ $0.34$ $0.14$
-npop2-cpop2-fb95-nTF-nodyn-NF $0.12$ $0.29$ $0.41$ $0.14$ $0.04$
-666 Table presents the fractional distribution of ULXs across different groups for various simulations. For group definitions see Table 3
+Donors of types MS, HG, and RG are all well-represented, suggesting that ULXs can form and persist across a wide range of stellar evolutionary stages.
 Huge majority are ULXs with BH accretors (typically $>80\%$) formed early in cluster evolution.
 Table 8 shows a modest anticorrelation between the relative size of the $G_{\textrm{IMBH}}$ and BH ULXs (Kendall [^8] Kendall’s $\tau$ is preferable to Pearson’s $r$ when working with small samples and non-normal distributions, while still providing a similarly interpretable measure of association between $-1$ and $+1$.’s $\tau=-0.49,-0.38,-0.37$, for $G_{\textrm{BH,MS}}$, $G_{\textrm{BH,HG}}$, and $G_{\textrm{BH,CHeB}}$, respectively).
-The presence of IMBH results in a continuous formation of IMBH powered ULX though the cluster history which, being an independent sourse to other BH ULX groups present mostly in the early cluster evolution phases, lowers their fractional imput to the ULX population.
-IMBH quickly removes BHs and possibly NSs from the system (e.g.
-Hong et al.
-2020).
-3.4.1 IMBH ULXs
-In this study, we define intermediate-mass black holes (IMBHs) as those with masses exceeding $500\mathrel{\mathrm{M}_{\odot}}$.
+The presence of IMBH results in a continuous formation of IMBH powered ULX though the cluster history which, being an independent source to other BH ULX groups present mostly in the early cluster evolution phases, lowers their fractional input to the ULX population.
+IMBH quickly removes BHs and possibly NSs from the system (e.g., Hong et al., 2020).
+3.5 IMBH ULXs
+In this study, we define IMBHs as those with masses exceeding $500\mathrel{\mathrm{M}_{\odot}}$.
 IMBH ULXs were observed in 6 simulations, all featuring multiple stellar populations (npop2) and non-tidally filling initial conditions (nTF).
 When present, IMBHs contribute significantly to the ULX population, accounting for 16-44% of ULXs over a Hubble time (Table 8).
-IMBHs form rapidly, within several Myr after the zero-age main sequence (ZAMS), classifying them as ”fast IMBHs” (Giersz et al.
-2015).
+IMBHs form rapidly, within several Myr after the ZAMS, classifying them as ”fast IMBHs” (Giersz et al., 2015).
 In simulations with delayed second populations, IMBH formation occurs 100+ Myr after the initial population.
@@
 Table 3: IMBH ULX properties
-label $\mathrm{t}_{\mathrm{phys,min}}$ $\mathrm{t}_{\mathrm{phys,max}}$ duty-cycle [%] companions
-npop2-cpop05-fb10-nTF $7.95$ $13,306.24$ $0.06$ MS: 22, HeWD: 4, COWD: 3, TPAGB: 2, CHeB: 2
-npop2-cpop05-fb10-nTF-NF $100.63$ $14,925.72$ $0.02$ MS: 41, COWD: 5, RG: 3, HeWD: 2, CHeB: 1,
-EAGB: 1, TPAGB: 1, ONeWD: 1
-npop2-cpop05-fb95-nTF $11.49$ $13,791.13$ $0.87$ MS: 29, COWD: 13, HG: 1, HeWD: 1, RG: 1
-npop2-cpop05-fb95-nTF-NF $104.89$ $14,751.19$ $0.22$ MS: 89, COWD: 12, TPAGB: 2, HeWD: 2, CHeB: 2,
-HeGB: 1, ONeWD: 1, HG: 1
-npop2-cpop2-fb10-nTF $9.19$ $14,171.06$ $0.14$ MS: 32, COWD: 5, HeWD: 2, RG: 1, TPAGB: 1,
-ONeWD: 1
-npop2-cpop2-fb95-nTF $8.99$ $14,546.35$ $0.43$ MS: 43, COWD: 11, HG: 3, HeWD: 3, TPAGB: 2,
-RG: 2
-888 Table presents the properties of $6$ IMBH which appeared in our simulations.
+label $\mathrm{t}_{\mathrm{phys,min}}$ $\mathrm{t}_{\mathrm{phys,max}}$ $\text{M}_{\mathrm{min}}$ $\text{M}_{\mathrm{max}}$ duty-cycle [%] companions
+npop2-cpop05-fb10-nTF $0.008$ $13.3$ $3.8$ $13.0$ $0.06$ MS: 22, HeWD: 4, COWD: 3,
+TPAGB: 2, CHeB: 2
+npop2-cpop05-fb10-nTF-NF $0.1$ $14.9$ $0.5$ $13.5$ $0.02$ MS: 41, COWD: 5, RG: 3,
+HeWD: 2, CHeB: 1, EAGB: 1,
+TPAGB: 1, ONeWD: 1
+npop2-cpop05-fb95-nTF $0.01$ $13.8$ $5.0$ $13.5$ $0.87$ MS: 29, COWD: 13, HG: 1,
+HeWD: 1, RG: 1
+npop2-cpop05-fb95-nTF-NF $0.1$ $14.8$ $0.6$ $12.9$ $0.22$ MS: 89, COWD: 12, TPAGB: 2,
+HeWD: 2, CHeB: 2, HeGB: 1,
+ONeWD: 1, HG: 1
+npop2-cpop2-fb10-nTF $0.009$ $14.2$ $0.6$ $4.5$ $0.14$ MS: 32, COWD: 5, HeWD: 2,
+RG: 1, TPAGB: 1, ONeWD: 1
+npop2-cpop2-fb95-nTF $0.009$ $14.5$ $2.1$ $6.8$ $0.43$ MS: 43, COWD: 11, HG: 3,
+HeWD: 3, TPAGB: 2, RG: 2
+999 Properties of $6$ intermediate-mass black holes (IMBHs) ULXs observed in our simulations.
+Each IMBH forms multiple, short-lived ULX phases with different companions.
+Column definitions: label - simulation model name; $\mathrm{t}_{\mathrm{phys,min}}$ minimum ULX phase onset age ($\mathrel{\mathrm{Gyr}}$); $\mathrm{t}_{\mathrm{phys,max}}$ maximum age at which ULX phases occur ($\mathrel{\mathrm{Gyr}}$); $\text{M}_{\mathrm{min}}$ minimum IMBH mass during ULX phases ($1000\;\times\mathrel{\mathrm{M}_{\odot}}$); $\text{M}_{\mathrm{max}}$ maximum IMBH mass during ULX phases ($1000\;\times\mathrel{\mathrm{M}_{\odot}}$); duty-cycle - percentage of total IMBH lifetime spent in ULX phases; companions - types and numbers of donor stars (MS - main sequence, HG - Hertzsprung gap, RG - red giant, CHeB - core helium burning, EAGB - early asymptotic giant branch, TPAGB - thermally pulsating asymptotic giant branch, HeGB - helium giant branch, HeWD - helium white dwarf, COWD - carbon-oxygen white dwarf, ONeWD - oxygen-neon white dwarf).
 […]
 IMBH ULXs formed exclusively in simulations with two stellar populations and non-tidally filling initial conditions.
-The centrally concentrated second population facilitates stellar interactions and mergers, while nTF clusters retain more stars, particularly black holes, in central regions, extending the IMBH formation window.
+The centrally concentrated second population facilitates stellar interactions and mergers, while nTF clusters retain more stars, particularly BHs, in central regions, extending the IMBH formation window.
+IMBH ULXs typically exist throughout the entire GC life, with first ULX phases appearing as early as $\sim 8\mathrel{\mathrm{Myr}}$ in some models and continuing until the end of our simulations at $\sim 14\mathrel{\mathrm{Gyr}}$, indicating that IMBH ULXs can form very late in GC evolution.
 3.6 ULX progenitors
@@
 D $4.73^{+31.64}_{-4.61}$ $0.14$ $0.80$ A-D(p) (52%) $0.28$ $0.66$ $22.43^{+46.86}_{-22.31}$ $1409.40^{+338400.60}_{-1396.45}$ $0.31^{+0.59}_{-0.31}$
-999 Values are provided as median with 95% confidence intervals or averages.
+101010 Initial properties of ULX progenitors, including median values with $95\%$ confidence intervals or averages.
 Groups are defined in Table 2 and accompaning text.
@@
 npop2-cpop2-fb95-nTF-NF $17$ $60$ $124$ $130$ $45$
+Approximate Average $19$ $62$ $40$ $51$ $11$
 111111 Distribution of ULX descendant fates for various simulation configurations.
@@
 Table 5 summarizes the various outcomes for ULX systems across different simulation configurations.
-ULX systems can undergo several fates: disruption, escape from the cluster, exchange interactions, mergers, or other outcomes.
+The table shows, that ULX systems can undergo several fates: disruption, escape from the cluster, exchange interactions, mergers, or other outcomes.
 The distribution of these fates appears largely stochastic across different simulation parameters.
-Systems categorized as ”other” typically experience no significant events for the remainder of their evolution and are predominantly NS/BH + WD or DCOs at the end of simulation.
+Systems categorized as ”other” typically experience no significant events for the remainder of their evolution and are predominantly NS/BH + WD or double compact objects at the end of simulation.
 Merger events typically occur $0.79^{+0.74}_{-0.53}\mathrel{\mathrm{Myr}}$ after the ULX phase.
-Double compact objects (DCOs) form in $\sim 63\%$ of cases, with $\sim 34\%$ of them being ejected from the cluster before merger or the end of simulation.
-These DCOs that remain in the cluster may contribute to the population of gravitational wave (GW) sources with merger rate of $\sim 11\%$.
+Double compact objects form in $\sim 63\%$ of cases, with $\sim 34\%$ of them being ejected from the cluster before merger or the end of simulation.
+These double compact objects that remain in the cluster may contribute to the population of gravitational wave sources with merger rate of $\sim 11\%$.
 Disruptions may result from dynamical interactions or binary evolution.
@@
 They tend to have lower-mass accretors and donors, with median masses of $13.00$ and $6.05\mathrel{\mathrm{M}_{\odot}}$, respectively, compared to $22.24$ and $39.34\mathrel{\mathrm{M}_{\odot}}$ for in-cluster ULXs.
-Escaper ULXs also have tighter orbits (median $57.42$ vs.
-$119.93\mathrel{\mathrm{R}_{\odot}}$) and a significantly higher fraction of neutron star accretors ($40\%$ vs.
-$4\%$).
+Escaper ULXs also have a significantly higher fraction of NS accretors ($40\%$ vs.
+$4\%$) and tighter orbits (median $57.42$ vs.
+$119.93\mathrel{\mathrm{R}_{\odot}}$).
+This orbital difference largely stems from dynamical hardening processes that occur prior to ejection, whereby close encounters cause binaries to become more tightly bound (Heggie, 1975).
+Additionally, the ejection mechanisms preferentially remove harder binaries with sufficient kinetic energy to escape the cluster potential.
 The median maximum X-ray luminosity of escaper ULXs ($\log_{10}\mathrm{L}_{\mathrm{X,max}}=39.89\mathrel{\mathrm{erg}/\mathrm{s}}$) is slightly lower than that of in-cluster ULXs ($\log_{10}\mathrm{L}_{\mathrm{X,max}}=40.26\mathrel{\mathrm{erg}/\mathrm{s}}$).
@@
 The ejection processes preferentially remove less massive systems, explaining the lower masses observed in escaper ULXs.
-The higher fraction of neutron star accretors among escapers could be due to the higher natal kicks and retention of more massive black holes within the cluster potential.
+This mass-dependent retention occurs because more massive systems require stronger dynamical encounters to achieve escape velocity from the cluster potential.
+The higher fraction of NS accretors among escapers could be due to the higher natal kicks and retention of more massive BHs within the cluster potential.
 These findings have important implications for understanding the origin and evolution of field ULXs.
-A significant fraction of observed field ULXs may have originated in globular clusters, which could explain some of the observed diversity in ULX populations.
+A significant fraction of observed field ULXs may have originated in GCs, which could explain some of the observed diversity in ULX populations.
+While our current analysis captures the properties of escaper ULXs at the moment of escape and during the ULX phase, tracking their complete galactic trajectories would require additional post-processing to integrate their orbits through the galactic potential (e.g., Cabrera & Rodriguez, 2023).
 Furthermore, the differences between escaper and in-cluster ULXs provide valuable insights into how environment and dynamical history can influence the properties of these extreme systems.
@@
 The parameter choices were carefully tailored to conditions characteristic of GCs, informed by insights from analogous studies.
-A key limitation in the broader study of ULXs in GCs is the paucity of confirmed observations within the Milky Way and its immediate vicinity.
-The substantial distances involved often complicate the detection and characterization of low-luminosity donor stars, which are believed to comprise a significant fraction of ULX donors (e.g.
-Wiktorowicz et al.
-2017).
+A key limitation in the broader study of ULXs in GCs is the dearth of confirmed observations within the Milky Way and its immediate vicinity.
+The brightest Milky Way X-ray binaries fall well below the ULX luminosity threshold (Fortin et al., 2023, 2024).
+The substantial distances involved often complicate the detection and characterization of low-luminosity donor stars, which are believed to comprise a significant fraction of ULX donors (e.g., Wiktorowicz et al., 2017).
 Furthermore, a critical challenge for this study lies in the difficulty of unambiguously differentiating GC-associated ULXs from those in the field environment.
@@
 Our escaper NS ULXs have a median maximum X-ray luminosity of $\log_{10}\mathrm{L}_{\mathrm{X,max}}=39.22\mathrel{\mathrm{erg}/\mathrm{s}}$, with a few BH ULXs exceeding $10^{40}\mathrel{\mathrm{erg}/\mathrm{s}}$.
-This stays in contrast to the field populations, where a significant fraction of more luminous sources is also expected (e.g., Wiktorowicz et al.
-2015).
-In the case of globular clusters, we find that such extreme ULXs are more likely to be retained within the cluster.
+This stays in contrast to the field populations, where a significant fraction of more luminous sources is also expected (e.g., Wiktorowicz et al., 2015).
+In the case of GCs, we find that such extreme ULXs are more likely to be retained within the cluster.
 […]
@@
 4.2 Counterparts
-The identification and characterization of ULX counterparts is crucial for understanding the nature of these systems, yet it remains a significant observational challenge (Heida et al.
-2016; Wiktorowicz et al.
-2021).
-While many ULXs are found in star-forming regions, their faintness in the optical and near-infrared (NIR), combined with the crowded environments, makes it difficult to confirm donor stars (López et al.
-2020; Allak et al.
-2022; Heida et al.
-2019).
-The observed optical emission can be a combination of light from the accretion disc and/or the donor star (Heida et al.
-2016).
-Accurate astrometry is essential for identifying counterparts, particularly for distinguishing between potential candidates within the error radii of ULX positions (Allak et al.
-2022).
-Red supergiants (RSGs) have also been identified as potential donors in ULXs (Heida et al.
-2016, 2019; López et al.
-2020).
-However, several sources initially classified as RSGs have been found to be nebulae, making the classification challenging (López et al.
-2020).
-Actually, the fraction of ULXs with RSG donors is potentially higher than predicted by binary evolution models (López et al.
-2020).
-It’s important to note that observational biases exist that favour the detection of more luminous and massive companions, such as RSGs (e.g.
-Heida et al.
-2019).
-Additionally, systems may be old, which makes it less likely for the donor to be a supergiant (Wiktorowicz et al.
-2017).
-Therefore, ULX counterparts remain ’candidates’ and need more thorough survey searches (Wiktorowicz et al.
-2021).
+Identifying ULX counterparts is crucial yet challenging (Heida et al., 2016; Wiktorowicz et al., 2021).
+While field ULXs face issues of faintness and crowded environments (López et al., 2020; Allak et al., 2022; Heida et al., 2019), extragalactic GC ULXs present unique challenges due to their unresolved stellar populations.
+However, GCs’ well-understood stellar demographics suggest these ULXs are likely LMXBs.
+Rare cases allow indirect donor identification, such as oxygen emission indicating a WD donor in a NGC 4472 GC ULX (Steele et al., 2014).
+Accurate astrometry remains essential for all ULX studies (Allak et al., 2022), and most counterparts, in both field and GCs, remain tentative due to observational biases (Heida et al., 2019; Wiktorowicz et al., 2021).
 4.3 Expectations for Antennae galaxy
-The correlation between X-ray source positions and stellar clusters has been previously noted in starburst galaxies (e.g.
-Kaaret et al.
-2004).
-The study of the Antennae galaxies by Poutanen et al.
-(2013) reveals a highly significant association between bright X-ray sources and stellar clusters, although most of these X-ray sources are located outside of the clusters.
-This finding is consistent with the idea that many ULXs are massive X-ray binaries that have been ejected from their birth clusters, rather than being intermediate mass black holes (IMBHs).
-Both ULXs and XRBs show a strong association with young stellar clusters, indicating a common origin.
-The displacements of ULXs from centres of their neighbouring star clusters are statistically significant and range up to 300 parsecs.
-This is likely due to the ejection of massive binaries from the clusters.
-The ejection mechanism suggested by Poutanen et al.
-(2013), supernova kicks and few-body encounters are consistent with the results of our work, where ULX among escapers have much higher NS accretor fraction and much smaller separations.
-Poutanen et al. (2013) provides evidences that dynamical ejection can be a major formation mechanism for these systems in the field.
+The correlation between X-ray source positions and stellar clusters has been previously noted in starburst galaxies (e.g., Kaaret et al., 2004).
+Poutanen et al.
+(2013)’s study of the Antennae galaxies reveals a significant association between bright X-ray sources and stellar clusters, with most X-ray sources located outside clusters.
+This supports the idea that many ULXs are massive X-ray binaries ejected from their birth clusters, rather than IMBHs.
+The displacements of ULXs from cluster centers, ranging up to $300$ parsecs, are statistically significant and likely due to ejection of massive binaries.
+The ejection mechanisms suggested by Poutanen et al.
+(2013), including supernova kicks and few-body encounters, are consistent with our results showing higher NS accretor fractions and smaller separations among escaped ULXs.
+The distribution of ULXs likely evolves with stellar population age.
+While the Antennae study focus on young populations, older stellar systems are expected to decrease their X-ray binary populations (e.g., Fragos et al., 2008), highlighting the importance of considering population age in ULX distribution studies.
 4.4 Future prospects
-Future studies of ULXs in globular clusters should include more simulations to better estimate statistical effects, errors, and the significance of conclusions.
+Future studies of ULXs in GCs should include more simulations to better estimate statistical effects, errors, and the significance of conclusions.
 A more consistent survey of the parameter space is needed, as well as the incorporation of evolutionary parameters (both stellar and binary evolution) beyond dynamical ones.
-In the future work we plan to include the effect of beaming in ULXs (e.g.
-Lasota & King 2023).
-Beaming increases the apparent luminosity of ULXs, but decreases the probability of observation due to possible misalignment (Wiktorowicz et al.
-2019; Khan et al.
-2022).
+In the future work we plan to include the effect of beaming in ULXs (e.g., Lasota & King, 2023).
+Beaming increases the apparent luminosity of ULXs, but decreases the probability of observation due to possible misalignment (Wiktorowicz et al., 2019; Khan et al., 2022).
 Population synthesis simulations, Wiktorowicz et al.
-(e.g.
-2019) indicate that the majority of neutron star (NS) ULXs are beamed.
-The beaming factor is dependent on the mass transfer rate, with higher mass transfer rates resulting in stronger beaming (King 2009).
-Additionally, the importance of wind Roche-lobe overflow (WRLOF; see e.g., Wiktorowicz et al.
-2021; Zuo et al.
-2021, and references therein) for NS ULXs with (super)giant donors should be explored.
-WRLOF can lead to stable mass transfer even with large mass ratios, and can boost mass transfer rates to reach ULX luminosity levels.
+(e.g., 2019) indicate that the majority of NS ULXs are beamed.
+The beaming factor is dependent on the mass transfer rate, with higher mass transfer rates resulting in stronger beaming (King, 2009).
+Additionally, the importance of wind Roche-lobe overflow (see e.g., Wiktorowicz et al., 2021; Zuo et al., 2021, and references therein) for NS ULXs with (super)giant donors should be explored.
+Wind Roche-lobe overflow can lead to stable mass transfer even with large mass ratios, and can boost mass transfer rates to reach ULX luminosity levels.
 5 Conclusions
@@
 The apparent absence of ULXs in Milky Way GCs aligns more closely with our models of initially tidally filling clusters rather than non-tidally filling ones.
+This distinction is crucial: tidally filling clusters experience significant early mass loss through tidal stripping by their host galaxy’s gravitational field, leading to the ejection of potential ULX progenitors before they can evolve into ULXs.
+In contrast, non-tidally filling clusters maintain higher central densities for most of their life, allowing for enhanced ULX formation through dynamical interactions.
+The presence of IMBH ULXs exclusively in non-tidally filling simulations further emphasizes the importance of maintaining dense stellar environments for certain ULX formation channels.
 Furthermore, our results suggest that field populations may be significantly ”polluted” by ULXs ejected from GCs (escapers), which could contribute to the relative underrepresentation of ULXs in GCs compared to field environments.
-Escapers can have properties very similar to the field populations or form dinsinct configurations.
+Escapers can have properties very similar to the field populations or form distinct configurations.
 Future observational efforts should prioritize the comparison of ULX properties across diverse environments, including GCs and the galactic field.
 Such studies will be critical for validating our theoretical predictions and advancing our understanding of the diverse formation channels of ULXs.
-Acknowledgments
+Acknowledgements.
+We are grateful to anonymous referee for their insightful comments which helped to improve the paper.
 GW, MG, AH, LH were supported by the Polish National Science Center (NCN) through the grant 2021/41/B/ST9/01191 AA acknowledges support for this paper from project No.
@@
 Data availability
-Data available on request.
+The data underlying this article are available in Zenodo at https://doi.org/10.5281/zenodo.14953837.
+This includes simulation data on ULX systems, globular cluster properties, and initial binary conditions, along with corresponding header files containing column definitions.
+Appendix A Additional tables
+Table 7: Model parameters and initial properties
+Label $\mathrm{N}_{\mathrm{pop}}$ $\mathrm{R}_{\mathrm{h,2}}/\mathrm{R}_{\mathrm{h,1}}$ $\mathrm{f}_{\mathrm{bin}}$ $\mathrm{R}_{\mathrm{gc}}$ $\mathrm{R}_{\mathrm{h}}$ $\mathrm{M}_{\mathrm{tot}}/10^{6}$ $\mathrm{R}_{\mathrm{c}}$ $\mathrm{M}_{\mathrm{c}}/10^{3}$ $\log_{10}(\rho_{\mathrm{c}})$ $\mathrm{t}_{\mathrm{h,rel}}$
+npop1-fb10-TF 1 10% $2.01$ $11.90$ $0.69$ $5.85$ $89.10$ $2.64$ $10,888.00$
+npop1-fb10-nTF 1 10% $2.01$ $1.00$ $0.69$ $0.49$ $89.10$ $5.87$ $265.22$
+npop1-fb95-TF 1 95% $1.96$ $11.93$ $0.72$ $5.81$ $93.93$ $2.61$ $6,386.70$
+npop1-fb95-nTF 1 95% $1.96$ $1.00$ $0.72$ $0.49$ $93.93$ $5.84$ $155.06$
+npop2-cpop05-fb10-TF 2 0.05 10% $2.01$ $11.69$ $0.66$ $0.13$ $12.71$ $356$,505.306505.306,505.306 , 505.30
+npop2-cpop05-fb10-TF-NF 2 0.05 10% $2.40$ $11.81$ $0.46$ $6.26$ $68.94$ $2.41$ $9,157.70$
+npop2-cpop05-fb10-nTF 2 0.05 10% $2.01$ $0.99$ $0.66$ $0.01$ $12.71$ $9.57$ $160.33$
+npop2-cpop05-fb10-nTF-NF 2 0.05 10% $2.40$ $1.00$ $0.46$ $0.53$ $68.94$ $5.63$ $225.73$
+npop2-cpop05-fb95-TF 2 0.05 95% $1.96$ $11.75$ $0.69$ $0.12$ $13.03$ $6.37$ $3,790.60$
+npop2-cpop05-fb95-TF-NF 2 0.05 95% $2.35$ $11.77$ $0.48$ $6.18$ $71.69$ $2.39$ $5,342.70$
+npop2-cpop05-fb95-nTF 2 0.05 95% $1.96$ $1.00$ $0.69$ $0.01$ $13.03$ $9.59$ $93.85$
+npop2-cpop05-fb95-nTF-NF 2 0.05 95% $2.35$ $1.00$ $0.48$ $0.53$ $71.69$ $5.61$ $132.31$
+npop2-cpop2-fb10-TF 2 0.2 10% $2.01$ $11.80$ $0.66$ $0.50$ $12.84$ $4.55$ $7,083.50$
+npop2-cpop2-fb10-TF-NF 2 0.2 10% $2.40$ $11.81$ $0.46$ $6.09$ $64.82$ $2.40$ $9,157.10$
+npop2-cpop2-fb10-nTF 2 0.2 10% $2.01$ $1.00$ $0.66$ $0.04$ $12.84$ $7.77$ $174.58$
+npop2-cpop2-fb10-nTF-NF 2 0.2 10% $2.40$ $1.00$ $0.46$ $0.52$ $64.82$ $5.61$ $225.72$
+npop2-cpop2-fb95-TF 2 0.2 95% $1.96$ $11.77$ $0.69$ $0.49$ $13.12$ $564$,079.504079.504,079.504 , 079.50
+npop2-cpop2-fb95-TF-NF 2 0.2 95% $2.35$ $11.77$ $0.48$ $6.12$ $70.32$ $2.35$ $5,342.50$
+npop2-cpop2-fb95-nTF 2 0.2 95% $1.96$ $1.00$ $0.69$ $0.04$ $13.12$ $7.77$ $101.01$
+npop2-cpop2-fb95-nTF-NF 2 0.2 95% $2.35$ $1.00$ $0.48$ $0.52$ $70.32$ $5.56$ $132.31$
+131313 List of simulations with differentiating parameters and some initial properties (at $\mathrm{t}_{\mathrm{phys}}=0$).
+Label identifies particular simulation and will be used throughout the paper.
+TF/nTF in the label means initially tidally filling or non-tidally filling.
+NF, when present, means that new features were included in this model (see text for details); $\mathrm{N}_{\mathrm{pop}}$ number of stellar generations; $\mathrm{R}_{\mathrm{h,2}}/\mathrm{R}_{\mathrm{h,1}}$ ratio between the $\mathrm{R}_{\mathrm{h}}$ of the second and first stellar generation; $\mathrm{f}_{\mathrm{bin}}$ binary fraction; $\mathrm{R}_{\mathrm{gc}}$ galactocentric radius [$\mathrel{\mathrm{kpc}}$]; $\mathrm{R}_{\mathrm{h}}$ half-mass radius for the first population [$\mathrel{\mathrm{pc}}$]; $\mathrm{M}_{\mathrm{tot}}$ total stellar mass [$\mathrel{\mathrm{M}_{\odot}}$]; $\mathrm{R}_{\mathrm{c}}$ core radius [$\mathrel{\mathrm{pc}}$]; $\mathrm{M}_{\mathrm{c}}$ core mass [$\mathrel{\mathrm{M}_{\odot}}$]; $\rho_{\mathrm{c}}$ central density [$\mathrel{\mathrm{M}_{\odot}}/\mathrel{\mathrm{pc}}^{3}$]; $\mathrm{t}_{\mathrm{h,rel}}$ Spitzer half-mass relaxation time [$\mathrel{\mathrm{M}_{\odot}}$].
+All simulations include the escapers information (see text).
+For all these simulations the counterpart field population (with dynamics turned off) was calculated (nodyn in the label; not shown in the table) with the same parameters.
+Table 8: Fraction of ULXs in different groups across simulations
+label $G_{\textrm{BH,MS}}$ $G_{\textrm{BH,HG}}$ $G_{\textrm{BH,CHeB}}$ $G_{\textrm{IMBH}}$ $G_{\textrm{NS}}$ other
+npop1-fb10-TF $0.17$ $0.33$ $0.36$ $0.11$ $0.04$
+npop1-fb10-nTF $0.16$ $0.32$ $0.30$ $0.12$ $0.11$
+npop1-fb10-nTF-nodyn $0.18$ $0.36$ $0.37$ $0.05$ $0.04$
+npop1-fb95-TF $0.11$ $0.40$ $0.33$ $0.12$ $0.04$
+npop1-fb95-nTF $0.12$ $0.39$ $0.28$ $0.11$ $0.11$
+npop1-fb95-nTF-nodyn $0.11$ $0.43$ $0.31$ $0.11$ $0.04$
+npop2-cpop05-fb10-TF $0.09$ $0.39$ $0.29$ $0.09$ $0.13$
+npop2-cpop05-fb10-nTF $0.07$ $0.29$ $0.20$ $0.33$ $0.04$ $0.06$
+npop2-cpop05-fb10-nTF-nodyn $0.10$ $0.46$ $0.32$ $0.04$ $0.07$
+npop2-cpop05-fb10-TF-NF $0.15$ $0.38$ $0.27$ $0.05$ $0.14$
+npop2-cpop05-fb10-nTF-NF $0.06$ $0.25$ $0.16$ $0.44$ $0.05$ $0.03$
+npop2-cpop05-fb10-nTF-nodyn-NF $0.11$ $0.48$ $0.33$ $0.08$
+npop2-cpop05-fb95-TF $0.11$ $0.29$ $0.42$ $0.11$ $0.06$
+npop2-cpop05-fb95-nTF $0.11$ $0.25$ $0.35$ $0.16$ $0.07$ $0.06$
+npop2-cpop05-fb95-nTF-nodyn $0.12$ $0.29$ $0.41$ $0.14$ $0.04$
+npop2-cpop05-fb95-TF-NF $0.13$ $0.27$ $0.44$ $0.10$ $0.06$
+npop2-cpop05-fb95-nTF-NF $0.10$ $0.18$ $0.29$ $0.34$ $0.05$ $0.04$
+npop2-cpop05-fb95-nTF-nodyn-NF $0.12$ $0.30$ $0.43$ $0.12$ $0.04$
+npop2-cpop2-fb10-TF $0.10$ $0.41$ $0.30$ $0.09$ $0.10$
+npop2-cpop2-fb10-nTF $0.06$ $0.28$ $0.19$ $0.36$ $0.04$ $0.07$
+npop2-cpop2-fb10-nTF-nodyn $0.10$ $0.46$ $0.32$ $0.04$ $0.07$
+npop2-cpop2-fb10-TF-NF $0.12$ $0.39$ $0.30$ $0.09$ $0.09$
+npop2-cpop2-fb10-nTF-NF $0.16$ $0.16$ $0.09$ $0.41$ $0.18$
+npop2-cpop2-fb10-nTF-nodyn-NF $0.11$ $0.48$ $0.33$ $0.08$
+npop2-cpop2-fb95-TF $0.13$ $0.26$ $0.44$ $0.13$ $0.05$
+npop2-cpop2-fb95-nTF $0.10$ $0.24$ $0.31$ $0.20$ $0.09$ $0.06$
+npop2-cpop2-fb95-nTF-nodyn $0.12$ $0.29$ $0.42$ $0.13$ $0.04$
+npop2-cpop2-fb95-TF-NF $0.12$ $0.28$ $0.45$ $0.11$ $0.05$
+npop2-cpop2-fb95-nTF-NF $0.13$ $0.17$ $0.22$ $0.34$ $0.14$
+npop2-cpop2-fb95-nTF-nodyn-NF $0.12$ $0.29$ $0.41$ $0.14$ $0.04$
+141414 Fractional distribution of ULXs across different groups for various simulations. For group definitions see Table 2